### v1

# Turning English-centric LLMs Into Polyglots: How Much Multilinguality Is Needed?

Abstract
The vast majority of today’s large language models are English-centric, having been pretrained predominantly on English text. Yet, in order to meet user expectations, models need to be able to respond appropriately in multiple languages once deployed in downstream applications. Given limited exposure to other languages during pretraining, cross-lingual transfer is important for achieving decent performance in non-English settings. In this work, we investigate just how much multilinguality is required during finetuning to elicit strong cross-lingual generalisation across a range of tasks and target languages. We find that, compared to English-only finetuning, multilingual instruction tuning with as few as three languages significantly improves a model’s cross-lingual transfer abilities on generative tasks that assume input/output language agreement, while being of less importance for highly structured tasks. Our code and data is available at https://github.com/ZurichNLP/multilingual-instruction-tuning.
Turning English-centric LLMs Into Polyglots:
How Much Multilinguality Is Needed?
[…]
1 Introduction
Conversational instruction tuning is a popular method for aligning large pretrained language models (LLMs) with user expectations such that they can effectively respond to a user’s input query and follow natural language instructions Ouyang et al. (2022); Wang et al. (2023); Chiang et al. (2023). An implicit expectation of conversational chatbots is that the language of a model’s response should match that of the user’s input query. For instance, unless otherwise specified, a German-language input should result in a German-language output. However, since the vast majority of pretraining and tuning data used to develop these models is in English, many instruction-tuned LLMs struggle to respond consistently in other languages Touvron et al. (2023b); Chen et al. (2023); Ye et al. (2023); Zhang et al. (2023b).
Figure 1: Input/output (IO) language agreement for English (en), German (de), Bulgarian (bg) and Icelandic (is) when instruction tuning on monolingual English (Mono) or on multilingual data (Multi-Guanaco). Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Despite this limited exposure, English-centric LLMs such as Llama 2 can achieve near-perfect input/output (IO) language agreement when tuned with relatively few multilingual conversational instructions. Figure 1 depicts IO language agreement, as measured by OpenLID [^1] https://github.com/laurieburchell/open-lid-dataset Burchell et al. (2023) and compares multilingual tuning on the language-diverse Guanaco dataset Dettmers et al. (2023); Köpf et al. (2023) with monolingual tuning on an English-only subset of instructions. As can be seen, multilingual tuning elicits strong IO language agreement, well within the bounds of language identification error rates, for non-English languages seen to differing degrees during pretraining and finetuning without degrading performance on English.
This observation raises two major questions, which we aim to address in this paper: Q1: How much multilinguality is required during finetuning to elicit few- and even zero-shot cross-lingual generalisation in English-centric LLMs? Q2: Which languages and tasks benefit most from multilingual instruction tuning of English-centric LLMs?
To investigate these questions, we instruction-tune English-centric LLMs, varying the number of languages seen during finetuning and evaluate performance in multiple target languages and on different tasks. We consider both high- and low-resource language settings with different scripts and investigate performance on generative tasks, such as open-ended chat and extractive question answering, as well as structured tasks aimed at assessing commonsense reasoning and language understanding.
Our results indicate that multilingual instruction tuning is crucial for eliciting cross-lingual transfer on generative tasks that assume IO language agreement while being less important in structured tasks that are commonly used to benchmark LLM performance. Furthermore, we empirically show that only a small number of finetuning languages is required to promote cross-lingual transfer. This highlights that tuning data for all potential target languages is not necessary to derive a capable polyglot chat model from an English-centric LLM.
2 Related Work
2.1 Instruction Tuning LLMs
Instruction tuning describes a supervised finetuning (SFT) strategy that aims to provide a model with a diverse set of demonstrations of user input queries paired with desirable model outputs. Unlike task-specific SFT, instruction tuning aims to promote cross-task generalisation, allowing for a ‘generalist’ model that is capable of completing any text-based task on the basis of natural language instructions provided at inference time Mishra et al. (2022); Wei et al. (2022); Wang et al. (2022); Sanh et al. (2022); Longpre et al. (2023). Meanwhile, framing instructions in a conversational manner and over multiple dialogue turns has been shown to be effective at deriving performant chat models Taori et al. (2023); Conover et al. (2023); Chiang et al. (2023); Dettmers et al. (2023); Ding et al. (2023). Furthermore, LLM instruction tuning remains effective given relatively limited labelled data (Ouyang et al., 2022; Touvron et al., 2023b; Zhou et al., 2023), parameter efficient methods Hu et al. (2021); Zhang et al. (2023a) and model quantisation Dettmers et al. (2023); Li et al. (2023).
2.2 Cross-lingual Transfer in English-centric LLMs
Despite numerous pushes towards improving NLP support for non-English and low-resource languages Costa-jussà et al. (2022); Le Scao et al. (2023), the vast majority of today’s publicly available LLMs are English-centric. For instance, GPT-3’s training data consisted of approximately 93% English documents with the remaining 7% pertaining to other languages Brown et al. (2020). [^2] https://github.com/openai/gpt-3/blob/master/dataset_statistics This trend is further reflected in popular open-weight LLMs (see Table 1). One potential reason for this could be the “the curse of multilinguality” Conneau et al. (2020) which describes how as multilinguality increases, evaluation performance decreases. For example, multilingual LLMs have been shown to sacrifice performance on English tasks due to having to share a finite amount of pre-training tokens and model parameters across more languages Lin et al. (2022); Le Scao et al. (2022). Consequently, the cross-lingual transfer abilities of more performant, English-centric models is important for many NLP practitioners looking to deploy these models in real-world settings.
A number of works have reported on the multilingual abilities of proprietary models such as GPT-3 and its derivatives across a range of NLU and NLG benchmarks Lai et al. (2023a); Holmström et al. (2023); Armengol-Estapé et al. (2022). On tasks such as machine translation (MT), GPT-3.5-Turbo has been shown to outperform dedicated systems for some high-resource languages Hendy et al. (2023); Lu et al. (2023); Jiao et al. (2023); Bang et al. (2023), particularly when translating between two non-English languages Laskar et al. (2023). However, performance on low-resource languages generally lags behind.
Meanwhile, other studies have focused on open-weight LLMs. Ye et al. (2023) compare multilingual reasoning capabilities of pretrained BLOOM Le Scao et al. (2023) and LLaMA Touvron et al. (2023a) and find that despite minimal amounts of non-English pretraining data, LLaMA has strong cross-lingual transfer abilities. Muennighoff et al. (2023) find that multilingual multitask finetuning improves BLOOM’s zero-shot abilities across all target languages compared with English-only finetuning. Finally, Chen et al. (2023) show that large-scale multilingual instruction tuning can improve performance on open-ended chat in multiple target languages.
This last work is most closely related to our current work but differs in a few key factors. Firstly, we investigate multilingual chat capabilities of more recent English-centric models, pretrained with significantly more data. Secondly, inspired by Zhou et al. (2023) we adopt a less-is-more finetuning approach, controlling the amount and distribution of multilingual data. In doing so, we focus our analysis on the minimal amount of multilinguality needed to elicit cross-lingual transfer. Finally, our evaluations consider more target languages and distinct downstream tasks.
3 Experimental Setup
To explore the multilingual capabilities of English-centric LLMs, we instruction-tune a series of models on a fixed-size set of examples, varying the number of languages available. Following this, we evaluate the resulting models in multiple target languages on four distinct tasks that are representative of how LLMs may be used in downstream applications.
3.1 English-centric LLMs
A prevailing trend in the development of recent LLMs is a clear focus on scaling up the size of the pretraining corpus Hoffmann et al. (2022). For instance, open-weight LLMs such as Falcon Almazrouei et al. (2023) and Llama 2 Touvron et al. (2023b) were trained on 1.5 and 2 trillion tokens respectively. Yet, although these numbers far surpass the 300 billion tokens used to train GPT-3 Brown et al. (2020), the distribution of language coverage remains similar across models with more than 90% pertaining to English (see Table 1). For our main experiments, we focus on Llama 2 7b, but also consider Llama 2 70b (§5.2) and Falcon 7b (Appendix E) to study the effect of model scaling and different training approaches respectively.
3.2 Instruction-tuning Data
For instruction tuning, we take inspiration from Dettmers et al. (2023) and finetune on high-quality conversations from the OpenAssistant dataset Köpf et al. (2023). These conversations comprise multiple dialogue turns between crowdworkers who were asked to either interact with or assume the role of a helpful AI assistant.
In contrast to Dettmers et al. (2023), who use all 9,846 top-rated conversations to train their ‘Guanaco’ models, we subsample training instances from the Guanaco dataset in order to control the amount of multilinguality. Specifically, we sample 3,200 unique English instances as an initial monolingual dataset, which we refer to as ‘Mono’. To construct datasets for multilingual finetuning we sample 200 unique training examples from each of five most frequent non-English languages in Guanaco (Spanish, Russian, German, Chinese, French). Given these subsets, we incrementally substitute English examples in Mono for non-English ones, one language at a time, following the order of how frequently each language appears in Guanaco. The resulting multilingual datasets are denoted as Multi-$i$, where $i$ equals the number of distinct languages included. For comparison, we also train models on the full Guanaco dataset, which includes more than 30 distinct languages.
3.3 Instruction-tuning and Inference Settings
For efficient instruction tuning, we use LoRA adapters ($R=64$, $\alpha=16$), leveraging Hugging Face’s TRL library [^3] https://github.com/huggingface/trl. We train all models for 2,000 update steps, using an effective batch size of 64 and a learning rate of $1^{e-5}$. Sequences longer than 1024 tokens are truncated. For 7-billion parameter models, we use two NVIDIA GeForce RTX 3090 with 24GB of memory. The time required for each training run is approximately 8 hours. For the larger 70-billion parameter model, discussed in §5.2, we use the same hyperparameters and train on four NVIDIA A100 GPUs with 80GB of memory each. Here, a single training run takes approximately 20 hours.
At inference time, we load the instruction-tuned models with vLLM Kwon et al. (2023). For open-ended generation tasks, we use top-p sampling (p=0.9) with a temperature of 0.8 and frequency penalty of 0.2. For the more constrained generation tasks we reduce the temperature to 0.001. To account for randomness under these sampling parameters, we report aggregated results for all tasks by repeating inference experiments with three random seeds.
3.4 Evaluation Tasks and Languages
As evaluation tasks, we consider open-ended chat, extractive question answering, commonsense reasoning and natural language inference. As all of these tasks differ in terms of the availability and representation of ground-truth labels, we describe the specific evaluation strategies in the following section.
Finally, we investigate performance on different target languages, which we select based on the makeup of Llama 2’s pretraining data (see Table 1). In doing so, we aim to investigate i) the supervised setting, where a target language has been seen during pretraining and finetuning; ii) the zero-shot cross-lingual setting, where a target language may have been seen during pretraining but not during finetuning Wu and Dredze (2019); and iii) the extreme low-resource setting, where the target language has not intentionally been seen at any stage of the training procedure. According to Touvron et al. (2023b), German (de), French (fr), Swedish (sv), Chinese (zh), Spanish (es) and Russian (ru) represent the most frequently seen non-English pretraining languages with an estimated 3.4 to 2.6 billion tokens. Catalan (ca), Norwegian (no) and Bulgarian (bg) represent languages less frequently seen during pretraining with an estimated 800 to 400 million tokens. And finally, we select Icelandic (is), Hindi (hi) and Greek (el) to represent extremely low-resource languages, whose frequency in Llama 2’s pretraining data is not available but may still appear in small amounts due to intentional or unintentional contamination Blevins and Zettlemoyer (2022). Unfortunately, existing benchmark evaluation datasets do not cover all of these languages and thus we limit the target languages in those tasks to the available subset of our target languages.
4 Experiments
4.1 Chat
General-purpose chatbots are a popular application of instruction-tuned LLMs. To evaluate chat performance, we make use of the AlpacaEval prompt dataset Dubois et al. (2023), which includes a diverse set of prompts for open-ended questions, creative writing, brainstorming, and other tasks. We randomly sample 300 prompts and translate these into each target language. As a translation engine, we follow Lai et al. (2023b) and use GPT-3.5-Turbo [^4] To translate chat prompts from AlpacaEval, we use gpt-3.5-turbo-0613.. In contrast to using dedicated translation systems, employing GPT-3.5-Turbo for this purpose has the advantage of being able to explicitly specify instructions that allow for preserving code blocks, tables and terminology, which we include as part of our translation prompt (see Figure B). Furthermore, since GPT-3.5-Turbo is trained on instruction- and conversational-style data, we expect it to perform well at translating in this domain.
To automatically evaluate open-ended chat responses, we leverage LLM-as-a-judge (Zheng et al., 2023). Following Zhou et al. (2023), given an input prompt and model’s response, we ask GPT-3.5-Turbo [^5] For chat evaluation, we use gpt-3.5-turbo-1106 due to its longer context window and cheaper inference costs. to grade the helpfulness of the response on a 6-point Likert scale (see Figure 16 for the prompt used). For each evaluation instance, we provide the prompt and model-generated response directly in the target language, which we found to be on par with evaluating via first translating responses into English (cf. Hada et al., 2023) (see Appendix C for more details). As noted by Chen et al. (2023), GPT-3.5-Turbo sometimes ignores the fact that the output language differs from the input language. Therefore, we force a score of 1 (indicating least helpful) if the language of the response does not match the intended target language.
Results
Figure 2: Average helpfulness of chat responses from Llama 2 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 3: X-QuAD results for Llama 2 7b with incremental multilingual instruction tuning. Results are shown for both multilingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 2 shows the helpfulness scores assigned by GPT-3.5-Turbo for all models for each of the target languages.
For English, we observe a reasonably strong and stable performance across all incremental multilingual instruction tuning settings, indicating that chat performance does not diminish with the increase in non-English training examples. In contrast, for most non-English target languages, performance increases significantly when moving from monolingual to bilingual instruction tuning. Most notably though, performance tends to plateau when instruction tuning with as few as three different languages. This observation holds for languages in both the supervised setting and the zero-shot setting, indicating that three finetuning languages is sufficient to elicit zero-shot cross-lingual transfer on this task. For languages in the extremely low-resource category, performance remains low despite multilingual finetuning. Manual inspection of these outputs reveal that, while these responses generally look convincing at first glance and are sufficient for language identification (see Figure 1), they are mostly nonsensical. This suggests that useful zero-shot cross-lingual transfer abilities of Llama 2 7b are limited to languages seen with higher frequency during pretraining.
4.2 Extractive Question Answering
Figure 4: X-CSQA results for Llama 2 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 5: XNLI results for Llama 2 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
In contrast to open-ended questions commonly used to query LLMs in chat settings, extractive question answering requires the model to identify relevant answer spans within longer context passages provided as part of the prompt. This kind of task closely resembles a retrieval augmented generation (RAG) setting, which is a popular method for extending an LLM’s knowledge with additional data not available during training Lewis et al. (2020); Izacard and Grave (2021). To evaluate our models on this task in multiple target languages, we use XQuAD Artetxe et al. (2020). [^6] We report results measured on the validation split of XQuAD since labels for the test split are not publicly available. This provides 1,190 QA pairs that were professionally translated into different languages.
For this task, we consider both a multilingual prompt setting and a monolingual prompt setting. The former presents the task instruction in English and provides the context passage and question in the relevant target language (en:xx), while the latter is entirely in the target language (xx:xx). As a starting point, we borrow the English prompt from Lai et al. (2023a) and manually translate it into each of the target languages considered. Additionally, we include a standardised response prefix as part of the prompt, effectively force-decoding the response “Based on the passage, the answer to the question is”. This allows us to better isolate the relevant answer string in the generative model’s output. A response is considered correct if the model’s answer string matches the ground truth answer after minimal postprocessing. [^7] We find that some postprocessing of model outputs is necessary for certain languages. Specifically, when queried with German and Russian prompts, all models consistently repeated the question before providing the extracted answer. To handle such edge cases, we strip away the question and truncate the system output to a maximum of 50 characters or the first line break, whichever comes first. An example of our prompting strategy for this task and model outputs is shown in Table 2.
Results
Figure 3 shows the performance on each of our target languages in XQuAD given incremental multilingual instruction tuning.
Similar to the results on the chat task (§4.1), we observe no performance degradation on English as the ratio of multilingual instructions increases. For most non-English target languages in the supervised setting, we see that when using the multilingual prompting strategy (en:xx), performance remains relatively uniform, with the exception of Chinese. In contrast, given the monolingual prompting strategy (xx:xx), multilingual instruction tuning can lead to small but noticeable gains in performance over monolingual tuning. Specifically, for German and Chinese, performance improves when moving from monolingual to bilingual instruction tuning and again plateaus rather quickly with as few as three languages. For languages in the extremely low-resource setting (Hindi and Greek), multilingual instruction tuning fails to improve the model’s ability to handle these languages. While performance gains on this task are generally less pronounced than on the open-ended chat task, we note that zero-shot extractive QA is inherently challenging for LLMs tuned on conversational instructions as they tend to generate verbose responses rather than the single word or entity that correctly matches the ground truth.
4.3 Commonsense Reasoning
Effectively understanding natural language requires some representation of the natural world and how concepts can relate to one another. Commonsense knowledge describes the set of general facts that reflects this. To evaluate how well English-centric models can reason across languages, we use the X-CSQA dataset. [^8] We report results for X-CSQA measured on the validation set of 1,000 questions. This dataset consists of questions paired with multiple choice answers which aim at assessing general, language-agnostic world knowledge involving different types of commonsense reasoning Talmor et al. (2019); Lin et al. (2021). Given a question and a set of five possible answers from A-E, we prompt the model to output the letter corresponding to the answer that best answers the question. Again, we borrow the English prompt template for this task from Lai et al. (2023a) as a starting point and translate it into each target language. An example of the prompt used is given in Table 3. Finally we assess performance with the monolingual prompting strategy (xx:xx).
Results
Figure 4 shows the accuracy on X-CSQA given incremental multilingual instruction tuning. Again, for English, we observe that multilingual instruction tuning does not significantly degrade performance. However, for non-English target languages, incremental multilingual instruction tuning fails to deliver improvements over monolingual tuning. This contrasts with the results on the previous tasks considered.
Figure 6: Chat performance of Llama 2 7b given multilingual instruction tuning using native non-English (Native NE) examples compared to using translated non-English (MT NE) examples. Error bars show a confidence interval of 95%.
4.4 XNLI
Natural language inference (NLI) is an important skill for LLMs, especially as input and output sequences grow to consist of multiple sentences. Given two sentences, this task aims to recognise a relationship between them either as entailment, contradiction or neutral. To assess an LLM’s ability to solve this task given multilingual instruction tuning, we use XNLI (Conneau et al., 2018) and evaluate performance on the official test split.
For this task, we use the implementation in the lm-evaluation harness Gao et al. (2023) [^9] https://github.com/EleutherAI/lm-evaluation-harness, which uses language-specific prompts for each target language (i.e. monolingual prompting (xx:xx)). Instead of querying the model to generate the desired label, multiple queries are constructed for each test instance (one for each possible label) and scored by the model. The sequence with the highest likelihood is chosen as the model’s answer.
Results
Figure 5 shows the accuracy measured on our target languages from XNLI given incremental multilingual instruction tuning. Strikingly, performance for all target languages remains uniform regardless of the amount of multilinguality used in finetuning. Furthermore, the equal performance between the Guanaco baseline and our Mono and Multi-$i$ finetuned models suggests that scaling up instruction diversity and the number of languages seen during finetuning beyond just six languages is still insufficient for delivering noticeable gains on this task.
5 Further Analysis
The results on four distinct evaluation tasks show that multilingual instruction tuning benefits open-ended chat in non-English languages most strongly (§4.1). In this section, we focus on this task to investigate the impact of instruction diversity and model scaling.
5.1 Instruction Diversity vs. Language Diversity
Figure 7: Chat performance of Llama 2 70b (denoted by triangular points) and Llama 2 7b (semi-transparent bars) with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Diversity is a key factor for sample efficient instruction tuning Zhou et al. (2023). Since our experiments in §4 make use of native non-English training instances from the OpenAssistant dataset, a potential confounding factor could be that adding more languages during finetuning also introduces more diverse training instructions. To investigate this, we retrain Multi-$i$ models using translated instruction-tuning data from Mono in place of native non-English examples, following the same incremental recipe as described in §3.2. This ensures that the data distribution remains constant as multilinguality increases. As a translation engine, we again use GPT-3.5-Turbo [^10] Since conversational training instances can be quite long, sometimes exceeding the default 4k token context window, we use gpt-3.5-tubo-16k for this translation task. and the prompt template provided in Figure 10.
Figure 6 compares the resulting performance on the open-ended chat task on a subset of our target languages (results for the remaining target languages are provided in Figure 15). Notably, we observe no significant differences between tuning with distinctly native non-English examples compared to those derived via automatic translation. This indicates that the gains attributed to increased multilinguality are not conflated with an increase in the diversity of instructions.
5.2 Scaling up Model Size
In order to investigate the effect of model scaling, we also repeat our chat experiments using Llama 2 70b as the underlying base model. Figure 7 shows the resulting helpfulness scores assigned by the LLM judge. Most notably, performance on non-English languages seen relatively frequently during pretraining is dramatically improved, often matching that of English. Secondly, we observe that the larger model’s performance on most non-English target languages tends to plateau with just two finetuning languages, unlike the smaller model that required three. Finally, while the performance on languages in the extremely low-resource scenario remains low, it exhibits a substantial relative improvement compared to the 7-billion parameter model. These results indicate that model scaling is extremely beneficial for exploiting the multilingual capabilities in English-centric models. This aligns with the assertion of Shaham et al. (2023) that larger models are more adept at handling multilinguality.
6 Discussion
Our findings show that multilingual instruction tuning can elicit zero-shot cross-lingual transfer in English-centric models, though its effectiveness on downstream performance varies across tasks. Notably, we observe significant performance gains for open-ended chat and noticeable gains on extractive QA with monolingual prompting. In comparison, for highly structured tasks such as X-CSQA and XNLI that impose a strict constraint on the output space regardless of the input language (e.g., choosing from options ‘A’, ‘B’, ‘C’, etc.), multilingual instruction tuning does not seem to provide substantial benefits. This distinction highlights that multilingual instruction tuning is most beneficial for open-ended generative tasks where there is an implicit agreement between the input query language and the language output by the model.
We posit that our findings align with the superficial alignment hypothesis Zhou et al. (2023). While instruction tuning guides the model towards a desirable ‘subdistribution of formats’ to use, a small amount of multilingual instruction data allows the model to learn a simple mapping between input and output languages. In essence, it helps steer the model towards language-specific ‘subdistributions of formats’ that agree with the language of the input. Therefore, we conclude that multilingual instruction tuning is of most benefit in scenarios where IO language agreement is required.
Most surprisingly, we find that multilingual instruction tuning with just three languages is sufficient to encourage IO language agreement in smaller LLMs, resulting in improved cross-lingual transfer on generative tasks. As model size increases, however, the number languages required decreases from three to two, indicating that larger models can learn this mapping easier. In both model sizes, adding more instruction tuning languages – including the target language itself – provides no significant gains.
Despite benefits from cross-lingual transfer, our results highlight a clear gap in performance between English and non-English target languages on all tasks. This indicates that small-scale multilingual instruction tuning is insufficient to improve the intrinsic multilingual capabilities of English-centric LLMs and that more advanced interventions may be required (e.g. Pfeiffer et al., 2020; ImaniGooghari et al., 2023).
7 Conclusion
In this paper, we investigated the use of multilingual instruction tuning to elicit multilingual capabilities of English-centric LLMs. Our results show that finetuning with as few as three languages can promote cross-lingual transfer and allow models to better exploit the relatively small amounts of non-English data seen during pretraining. Experiments conducted on four distinct tasks revealed that this can lead to significant performance improvements on open-ended generative tasks that assume input and output language agreement. Furthermore, we observed that these effects are amplified in larger models, requiring only two finetuning languages for successful multilingual instruction tuning. The effectiveness of cross-lingual transfer, which reduces the need for extensive multilingual instruction tuning, is good news for LLM developers. However, future work could explore other methods to encourage smaller models to better leverage the limited non-English data seen during pretraining and whether there are tasks for which language-specific instruction tuning is of higher importance.
8 Limitations
This work focused on detecting and evaluating non-English responses from English-centric LLMs. Our experiments demonstrated that cross-lingual transfer, which can be elicited given small amount of multilingual instruction tuning, can be effective for improvinging the non-English performance of English-centric models. That said, our analysis focused primarily on Indo-European languages, except for Chinese. As a consequence, it remains unclear how well our results generalise to more distant or topologically diverse languages, however, performance gains observed on Chinese suggest that generalisation is more dependent on exposure during pretraining than language similarity.
For our experiments and analysis on model training data, we relied on automatic language identification. To this end, we employed the OpenLID model from Burchell et al. (2023). Despite low error rates achieved by this model, language identification is not perfect and can lead to some texts being misidentified. To mitigate the risk of language contamination Blevins and Zettlemoyer (2022) in our controlled experiments, we only included training examples whose language is identified with a confidence threshold $\geq$ 0.8.
In §5.1 we investigated the impact of multilingual diversity versus training example diversity. While our findings revealed that there is no significant difference between these two settings, we note that even when finetuning with the original native non-English examples, task diversity may be inherently limited by design of the data collection. For instance, regardless of the language used, crowdworkers were asked to follow the same set of guidelines [^11] https://projects.laion.ai/Open-Assistant/docs/guides/guidelines when creating the data.
Acknowledgements
We sincerely thank our friends and colleagues, Fabian Aiolfi, Thea Bächler, Anastassia Shaitarova, Finnur Ágúst Ingimundarson, Cui Ding, Andrianos Michail, Janine Aeberhard, Arnisa Fazla, Farhad Nooralahzadeh, Omnia Ibrahim, Xuan Lam and Manasi Muglikar, for helping us with language-specific questions and validating translations used in our experiments, as well as Lena Bolliger and Patrick Haller for providing valuable feedback. Rico Sennrich acknowledges funding by the Swiss National Science Foundation (project MUTAMUR; no. 213976).
Appendix A Pretraining Data for English-centric LLMs
LLMs are pretrained on enormous amounts of unlabelled text gathered from existing corpora and the web. With the notable exception of BLOOM (Le Scao et al., 2023), very limited information is offered about the distribution of languages represented in datasets used to train LLMs and the subsequent performance on non-English languages. Table 1 provides an overview of the document-level language distributions of the LLMs used in this paper. For Llama 2, the information is taken from the original paper Touvron et al. (2023b), in which the authors analyse the training data using a fastText language classifier on corpus documents with a threshold of 0.5. For GPT-3, we use the official dataset statistics made available on Github [^12] https://github.com/openai/gpt-3/blob/master/dataset_statistics/languages_by_document_count.csv which provide document-level language identification information. To gather statistics for Falcon, we inspected a sample of the RefinedWeb dataset Penedo et al. (2023a) that was constructed to train this family of models. [^13] https://huggingface.co/datasets/tiiuae/falcon-refinedweb Using the OpenLID fastText model from Burchell et al. (2023), we identify the most frequent document-level languages based on approximately 320 million examples from this corpus. Languages are identified with a confidence threshold of $\geq$ 0.6 and predictions below this threshold are aggregated under ‘unknown’.
Llama 2 Falcon GPT-3
# Language Percent # Language Percent # Language Percent
1 English 89.70% 1 English 94.790% 1 English 93.69%
2 unknown 8.38% 2 unknown 4.944% 2 German 1.20%
3 German 0.17% 3 Standard Malay 0.031% 3 French 1.02%
4 French 0.16% 4 Swahili 0.027% 4 Portuguese 0.64%
5 Swedish 0.15% 5 French 0.016% 5 Italian 0.58%
6 Chinese 0.13% 6 Portuguese 0.012% 6 Spanish 0.51%
7 Spanish 0.13% 7 Chinese (Traditional) 0.011% 7 Dutch 0.37%
8 Russian 0.13% 8 Sardinian 0.011% 8 Polish 0.25%
9 Dutch 0.12% 9 Dutch 0.009% 9 Japanese 0.25%
10 Italian 0.11% 10 Spanish 0.009% 10 Danish 0.16%
11 Japanese 0.10% 11 Italian 0.008% 11 Norwegian 0.15%
12 Polish 0.09% 12 German 0.008% 12 Romanian 0.13%
13 Portuguese 0.09% 13 Danish 0.005% 13 Finnish 0.13%
14 Vietnamese 0.08% 14 Russian 0.005% 14 Chinese (Simplified) 0.12%
15 Ukranian 0.07% 15 Indonesian 0.005% 15 Russian 0.11%
16 Korean 0.06% 16 Swedish 0.005% 16 Czech 0.10%
17 Catalan 0.04% 17 Somali 0.005% 17 Swedish 0.06%
18 Serbian 0.04% 18 Vietnamese 0.004% 18 Hungarian 0.06%
19 Bahasa Indonesian 0.03% 19 Venetian 0.004% 19 Chinese (Traditional) 0.04%
20 Czech 0.03% 20 Chinese (Simplified) 0.004% 20 Bahasa Indonesian 0.04%
21 Finnish 0.03% 21 Tagalog 0.004% 21 Croation 0.04%
22 Hungarian 0.03% 22 Northern Uzbek 0.003% 22 Turkish 0.04%
23 Norwegian 0.03% 23 Limburgish 0.003% 23 Catalan 0.03%
24 Romanian 0.03% 24 Polish 0.003% 24 Vietnamese 0.03%
25 Bulgarian 0.02% 25 Turkish 0.003% 25 Slovenian 0.03%
26 Danish 0.02% 26 Norwegian Nynorsk 0.003% 26 Estonian 0.02%
27 Slovenian 0.01% 27 Norwegian Bokmal 0.003% 27 Slovak 0.02%
28 Croatian 0.01% 28 Romanian 0.002% 28 Korean 0.02%
n/a Icelandic n/a 30 Greek 0.002% 31 Greek 0.017%
n/a Hindi n/a 60 Bulgarian 0.001% 45 Hindi 0.004%
n/a Greek n/a 72 Icelandic 0.001% 47 Icelandic 0.004%
48 Bulgarian 0.003%
102 Hindi 0.0001%
Table 1: Distribution of document languages in pretraining datasets for popular English-centric LLMs. For Llama 2 and GPT-3, statistics are taken from the original works. For Falcon, statistics are estimated based on a sample of the RefinedWeb corpus (roughly 320 million documents). Note that for Llama 2 and Falcon, ‘unknown’ indicates texts identified below a predefined confidence threshold (0.5 for Llama 2, 0.65 for our analysis with RefinedWeb). Touvron et al. (2023b) state that this data partially pertain to programming code. Target languages for our experiments are highlighted in bold.
Appendix B Translating AlpacaEval Prompts
You are a helpful assistant. Translate the following text into {{target_language}}. Keep the structure of the original text and preserve things like code and names. Please ensure that your response contains only the translated text. The translation must convey the same meaning as the original and be natural for native speakers with correct grammar and proper word choices. Your translation must also use exact terminology to provide accurate information even for the experts in the related fields. Original: {{source_text}} Translation into {{target_language}}:
Figure 8: Prompt template used to translate AlpacaEval with GPT-3.5-Turbo. Values in curly braces represent placeholders. The value specified for ‘target language’ is English name of the language (e.g., ‘German’) we are translating into, except for Greek and Chinese, where we specify standard modern Greek and Mandarin Chinese respectively.
To translate AlpacaEval prompts from English into each target language, we query GPT-3.5-Turbo using the template in Figure 8. To validate the translated prompts, we manually inspected a sample of the outputs in various languages. This inspection revealed that the translations were typically decent, although often included literal translations for metaphorical expressions rather than how a native speaker might express themselves. For instance, English ‘bullet points’ was translated literally into Hindi, rather than an arguably more appropriate phrasing such as ‘important points’. Additionally, in languages that distinguish between formal and informal or gendered pronouns (e.g., German, French, Hindi), the formal and male forms are dominant, which may be less representative of how native speakers actually interact with LLM chatbots.
Appendix C Direct vs. Translated Evaluation for Non-English Chat Responses
Figure 9: Distribution of helpfulness scores assigned by our LLM judge, GPT-3.5-Turbo, using direct evaluation in the specified target language and after translating target-language responses to English.
While using a powerful LLM to evaluate the outputs of other models has been shown to achieve reasonable agreement with human judgements in English (Zheng et al., 2023; Chiang and Lee, 2023), it is unclear whether this agreement transfers to all languages under investigation. Recent work by Hada et al. (2023) has shown that agreement between human and LLM judges tends to be lower for non-English languages, especially in the case of low-resource and non-Latin scripted languages, where the LLM judge tends to be overly optimistic in its assessment. However, for certain assessment criteria, such as linguistic acceptability and general content quality, they also confirm that inter-annotator agreement between LLM-based evaluators and humans is inline with that of multiple human annotators.
To investigate this potential bias, we compared scores assigned by the LLM judge on model-generated responses directly in each non-English target language and on their English translations. For each non-English prompt-response pair, we use GPT-3.5-Turbo (gpt-3.5-turbo-1106) and the prompt shown in Figure 8, specifying English as the target language. The resulting translated response is then paired with its corresponding AlpacaEval prompt in its original English form and presented to the LLM judge using the prompt in Figure 16.
Figure 9 shows that the distribution of assigned scores in the direct and translated evaluation settings is very similar for most languages. For languages that use non-Latin scripts (e.g., Chinese, Russian, Bulgarian), we observe that GPT-3.5-Turbo tends to assign slightly higher scores more frequently when evaluating directly on the non-English prompt/response pairs. This finding agrees with those from Hada et al. (2023) and indicates that LLM-based evaluations in non-English languages tend to be slightly overly optimistic and should be considered with caution. Nevertheless, we observe that the discrepancy between direct and translated evaluation is relatively minor for the languages considered and leads to negligible differences in the overall average helpfulness score. Because of this, and to keep evaluation costs to a minimum, we opt to use the direct evaluation strategy for our experiments.
Appendix D Translating Non-English Guanaco Training Examples into English
You are a helpful assistant. Translate the following conversation between a human and an AI assistant into {{target_language}}. Keep the structure of the original text and preserve things like code, names and role labels (e.g. <S1>, <S2>). Please ensure that your response contains only the translated text. The translation must convey the same meaning as the original and be natural for native speakers with correct grammar and proper word choices. Your translation must also use exact terminology to provide accurate information even for the experts in the related fields. Original: {{training_instance}} Translation into {{target_language}}
Figure 10: Prompt template used to translate Guanaco training instances from English into various target languages.
In order to investigate the effect of language diversity compared to instruction diversity, we translate a subset of Guanaco’s English training examples into non-English target languages and use these to create MT-based Multi-$i$ instruction-tuning datasets. By default, speaker roles in Guanaco are denoted with ‘### Human:’ and ‘### Assistant’. To ensure that these are never translated and the dialogue structure is maintained, we substitute them with special tokens ‘<S1>’ and ‘<S2>’ and explicitly tell the model to leave these tokens intact (see Figure 10). Before training, the special tokens back to their original form.
Appendix E Results with Falcon 7b
In order to assess whether our findings generalise to other LLMs, we repeat our experiments using Falcon 7b Almazrouei et al. (2023).
E.1 Chat
Figure 11 shows the helpfulness scores assigned by our LLM judge for Falcon 7b given incremental multilingual instruction tuning across all target languages. Similarly to our results with Llama 2 7b (see Figure 2), cross-lingual transfer is elicited after finetuning with relatively few languages, and no additional gains observed when including more than three languages. That said, Falcon 7b appears to show strong performance on French, even without multilingual finetuning, indicating that, despite being an English-centric model, it has strong capabilities in French out of the box. For Spanish, German and Chinese, performance is comparable to that of Llama 2 7b. However, for all other languages, responses are often ranked least helpful, indicating that Falcon 7b’s multilingual capabilities are limited strictly to major European languages and Chinese.
E.2 Extractive Question Answering
Figure 12 shows the results of Falcon 7b on XQuAD. While performance is generally lower than that of Llama 2 7b on this task (see Figure 3), we observe a similar, albeit weaker, effect of multilingual finetuning within the supported languages (Spanish, German, Chinese).
E.3 Commonsense Reasoning
Figure 13 shows the results of Falcon 7b on X-CSQA. Strikingly, in contrast to the results achieved with Llama 2 7b (Figure 4), Falcon 7b fails to score above random performance across all target languages. Regarding the effect of multilingual instruction tuning, we again see that it fails to deliver any performance improvements on this highly structured task.
E.4 XNLI
Figure 14 shows the results of Falcon 7b on XNLI. Similar to the results attained with Llama 2 7b (see Figure 5), we observe no significant differences in performance given different degrees of multilingual instruction tuning.
E.5 Discussion
Overall, these experiments indicate that Falcon’s multilingual capabilities are considerably narrower than Llama 2’s. We hypothesise that this may be a result of more stringent filtering of web-scraped pretraining data Penedo et al. (2023b), possibly resulting in less accidental contamination. Interestingly though, according to our estimated statistics based on the sample of the RefinedWeb corpus (Table 1), which makes up a large portion of Falcon’s pretraining data, Russian appears as frequently as Chinese and Swedish. However, we find that Russian texts are, on average, three times longer when tokenized with Falcon’s tokenizer compared to Llama 2’s. This indicates that, despite having a much larger vocabulary (65k for Falcon vs. 32k for Llama 2), Cyrillic-scripted languages such as Russian and Bulgarian are outside of the model’s intended domain and performance on these languages is inherently limited.
Figure 11: Average helpfulness of chat responses from Falcon 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 12: X-QuAD results for Falcon 7b with incremental multilingual instruction tuning. Results are shown for both multilingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 13: X-CSQA results for Falcon 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 14: XNLI results for Falcon 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 15: Comparison of Llama 2 7b with multilingual instruction tuning using native (Native NE) and translated non-English (MT NE) examples for additional target languages not shown in Figure 6. Error bars show a confidence interval of 95%.
You are an expert language evaluator. You are evaluating a response that has been submitted for a particular task, using a specific set of standards. Below is the data: [BEGIN DATA] *** [Task]: {{task}} *** [Submission]: {{submission}} *** [Criterion]: helpfulness: ‘‘1’’: ‘‘Not helpful - The generated text is completely irrelevant, unclear, or incomplete. It does not provide any useful information to the user.’’ ‘‘2’’: ‘‘Somewhat helpful - The generated text has some relevance to the user’s question, but it may be unclear or incomplete. It provides only partial information, or the information provided may not be useful for the user’s needs.’’ ‘‘3’’: ‘‘Moderately helpful - The generated text is relevant to the user’s question, and it provides a clear and complete answer. However, it may lack detail or explanation that would be helpful for the user.’’ ‘‘4’’: ‘‘Helpful - The generated text is quite relevant to the user’s question, and it provides a clear, complete, and detailed answer. It offers additional information or explanations that are useful for the user. However, some of the points of the response are somewhat repetitive or could be combined for greater clarity and concision.’’ ‘‘5’’: ‘‘Very helpful - The generated text is highly relevant to the user’s question, and it provides a clear, complete, and detailed answer. It offers additional information, explanations, or analogies that are not only useful but also insightful and valuable to the user. However, the structured of the response is not well-organized and there is no clear progression or logical sequence of different points in the response.’’ ‘‘6’’: ‘‘Highly helpful - The generated text provides a clear, complete, and detailed answer. It offers additional information or explanations that are not only useful but also insightful and valuable to the user. The response is also in a logical and easy-to-follow manner by explicitly using headings, bullet points, or numbered lists to break up the information and make it easier to read.’’ *** [END DATA] Does the submission meet the criterion? First, write out in a step by step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Provide your response as a RFC8259 compliant JSON following this schema: {‘‘reasoning’’: str, ‘‘score’’: int}
Figure 16: Prompt template used to query the LLM judge in order to assess the helpfulness of an open-ended chat response. This prompt closely follows the one used by Zhou et al. (2023) but extends it by specifying the return value as a valid JSON-like object.
Prompt:
### Human: Answer the question from the given passage. Your answer should be directly extracted from the passage, and it should be a single entity, name, or number, not a sentence.
Passage:
[…]
Question:
Mit wie vielen Interceptions erzielte Josh Norman im Jahr 2015 Touchdowns?
Note: Your answer should be directly extracted from the passage and be a single entity, name, or number. Provide the answer in quotations.
### Assistant: Based on the passage, the answer to the question is ‘‘
Reference:
zwei
System:
two".
In the passage, it is stated that Josh Norman “erzielte vier Interceptions, von denen zwei zu Touchdowns für sein Team wurden.” This means that Norman intercepted four passes during the season, and two of those interceptions were returned for touchdowns.
Table 2: Example input prompt, reference and model response for the XQuAD task. Monospaced font represents the prompt template used for each test instance. Here, we show the multilingual prompting strategy (en:xx), where the task instruction and response prefix is provided English, while the context passage and question are in the target language (in this case, German). A correct answer in this setting requires the model to continue the response by switching into the target language, which it fails to do in this particular example.
Prompt:
### Human: Dir wird eine Frage vorgelegt, die mehrere Antwortmöglichkeiten hat. Wähle die am besten geeignete Option aus ‘‘A’’, ‘‘B’’, ‘‘C’’, ‘‘D’’ oder ‘‘E’’ basierend auf deinem gesunden Menschenverstand. Bitte gebe deine Antwort in Form eines einzelnen Buchstabens in Anführungszeichen an.
Frage:
Sarah stellte ihr Essen in einen Mülleimer neben ihrem Schreibtisch. Wo mag sie gewesen sein?
Antwortmöglichkeiten:
A: Bushaltestelle
B: Ecke
C: Motel
D: Park
E: Büro
### Assistant: ‘‘
Reference:
E
System:
E”
Table 3: Example of the X-CSQA task. Monospaced font represents the prompt template used for each test instance. Here, the prompt strategy is monolingual (xx:xx) with the task instruction provided in the same language as the test instance (in this case, German). A correct answer requires the model to output the letter corresponding to the ground-truth answer.

### v2

@@
 Abstract
-The vast majority of today’s large language models are English-centric, having been pretrained predominantly on English text.
+The vast majority of today’s large language models (LLMs) are English-centric, having been pretrained predominantly on English text.
 Yet, in order to meet user expectations, models need to be able to respond appropriately in multiple languages once deployed in downstream applications.
-Given limited exposure to other languages during pretraining, cross-lingual transfer is important for achieving decent performance in non-English settings.
-In this work, we investigate just how much multilinguality is required during finetuning to elicit strong cross-lingual generalisation across a range of tasks and target languages.
-We find that, compared to English-only finetuning, multilingual instruction tuning with as few as three languages significantly improves a model’s cross-lingual transfer abilities on generative tasks that assume input/output language agreement, while being of less importance for highly structured tasks.
-Our code and data is available at https://github.com/ZurichNLP/multilingual-instruction-tuning.
+This requires strong cross-lingual transfer abilities.
+In this work, we investigate the minimal amount of multilinguality required during finetuning to elicit cross-lingual generalisation in English-centric LLMs.
+In experiments across four LLMs, we find that multilingual instruction tuning with as few as two to three languages is both necessary and sufficient to elicit effective cross-lingual generalisation, with the limiting factor being the degree to which a target language is seen during pretraining.
+Evaluations on five different tasks further reveal that multilingual instruction tuning is most beneficial for generative tasks that assume input/output language agreement, such as in chat settings, while being of less importance for highly structured classification-style tasks.
+[^1] Our code and data is available at https://github.com/ZurichNLP/multilingual-instruction-tuning.
 Turning English-centric LLMs Into Polyglots:
 How Much Multilinguality Is Needed?
-[…]
+Tannon Kew1thanks:   Work partially conducted during a research internship at Textshuttle.
+Florian Schottmann2,3       Rico Sennrich1,4 1University of Zurich, 2Textshuttle, 3ETH Zurich, 4University of Edinburgh {kew,sennrich}@cl.uzh.ch, schottmann@textshuttle.com
 1 Introduction
@@
 (2023); Chiang et al.
+(2023); Zhou et al.
 (2023).
@@
 For instance, unless otherwise specified, a German-language input should result in a German-language output.
-However, since the vast majority of pretraining and tuning data used to develop these models is in English, many instruction-tuned LLMs struggle to respond consistently in other languages Touvron et al.
-(2023b); Chen et al.
+However, since the vast majority of training data is in English, many instruction-tuned LLMs struggle to respond consistently in other languages Ouyang et al.
+(2022); Touvron et al.
+(2023); Chen et al.
 (2024); Ye et al.
@@
 (2024b).
-Figure 1: Input/output (IO) language agreement for English (en), German (de), Bulgarian (bg) and Icelandic (is) when instruction tuning on monolingual English (Mono) or on multilingual data (Multi-Guanaco).
-Striped bars indicate that the target language is included in the finetuning data.
+Figure 1: Input/output (IO) language agreement for English (en), German (de), Bulgarian (bg), and Icelandic (is) given English-only instruction tuningono) or andmultilingual instruction tuning (Multi-Guanaco).
+Striped bars indicate that the target language is not seen during finetuning (i.e.
+the 0-shot setting).
 Error bars show a confidence interval of 95%.
-Despite this limited exposure, English-centric LLMs such as Llama 2 can achieve near-perfect input/output (IO) language agreement when tuned with relatively few multilingual conversational instructions.
-Figure 1 depicts IO language agreement, as measured by OpenLID [^1] https://github.com/laurieburchell/open-lid-dataset Burchell et al.
-(2023) and compares multilingual tuning on the language-diverse Guanaco dataset Dettmers et al.
+Despite limited exposure, however, English-centric LLMs such as Llama 2 seemingly achieve near-perfect input/output (IO) language agreement when tuned with relatively few multilingual conversational instructions.
+Figure 1 depicts IO language agreement, as measured by OpenLID Burchell et al.
+(2023), and compares multilingual tuning on the language-diverse Guanaco dataset Dettmers et al.
 (2023); Köpf et al.
 (2023) with monolingual tuning on an English-only subset of instructions.
-As can be seen, multilingual tuning elicits strong IO language agreement, well within the bounds of language identification error rates, for non-English languages seen to differing degrees during pretraining and finetuning without degrading performance on English.
-This observation raises two major questions, which we aim to address in this paper: Q1: How much multilinguality is required during finetuning to elicit few- and even zero-shot cross-lingual generalisation in English-centric LLMs?
+As can be seen, multilingual tuning elicits strong IO language agreement – within the bounds of OpenLID’s error rates – for non-English languages seen to varying degrees during pretraining and finetuning without degrading performance on English.
+This observation raises two major questions, which we aim to address in this paper: Q1: How much multilinguality is required during finetuning to elicit 0-shot cross-lingual generalisation in English-centric LLMs?
 Q2: Which languages and tasks benefit most from multilingual instruction tuning of English-centric LLMs?
-To investigate these questions, we instruction-tune English-centric LLMs, varying the number of languages seen during finetuning and evaluate performance in multiple target languages and on different tasks.
-We consider both high- and low-resource language settings with different scripts and investigate performance on generative tasks, such as open-ended chat and extractive question answering, as well as structured tasks aimed at assessing commonsense reasoning and language understanding.
+To investigate these questions, we instruction tune four distinct English-centric LLMs with varying degrees of multilinguality and evaluate performance on a diverse set of up to 19 languages across five different tasks.
+Specifically, we consider high-, medium-, and low-resource target languages with regard to a model’s pretraining data distribution and generative tasks, such as single-turn dialogue, sentence simplification, and extractive question answering, as well as more structured tasks aimed at assessing commonsense reasoning and language understanding.
 Our results indicate that multilingual instruction tuning is crucial for eliciting cross-lingual transfer on generative tasks that assume IO language agreement while being less important in structured tasks that are commonly used to benchmark LLM performance.
 Furthermore, we empirically show that only a small number of finetuning languages is required to promote cross-lingual transfer.
-This highlights that tuning data for all potential target languages is not necessary to derive a capable polyglot chat model from an English-centric LLM.
+This highlights that tuning data for all potential target languages is not necessary to derive capable polyglot chat models from English-centric LLMs.
 2 Related Work
 2.1 Instruction Tuning LLMs
-Instruction tuning describes a supervised finetuning (SFT) strategy that aims to provide a model with a diverse set of demonstrations of user input queries paired with desirable model outputs.
-Unlike task-specific SFT, instruction tuning aims to promote cross-task generalisation, allowing for a ‘generalist’ model that is capable of completing any text-based task on the basis of natural language instructions provided at inference time Mishra et al.
+Unlike task-specific finetuning, instruction tuning aims to promote cross-task generalisation, allowing for a ‘generalist’ model that is capable of completing any text-based task on the basis of natural language instructions provided at inference time Mishra et al.
 (2022); Wei et al.
@@
 (2023).
-Furthermore, LLM instruction tuning remains effective given relatively limited labelled data (Ouyang et al., 2022; Touvron et al., 2023b; Zhou et al., 2023), parameter efficient methods Hu et al.
+Furthermore, LLM instruction tuning remains effective given relatively limited labelled data Ouyang et al.
+(2022); Touvron et al.
+(2023); Zhou et al.
+(2023), parameter efficient training strategies Hu et al.
 (2021); Zhang et al.
@@
 (2023); Li et al.
-(2023).
+(2023b).
 2.2 Cross-lingual Transfer in English-centric LLMs
-Despite numerous pushes towards improving NLP support for non-English and low-resource languages Costa-jussà et al.
-(2022); Le Scao et al.
-(2023), the vast majority of today’s publicly available LLMs are English-centric.
+The vast majority of today’s publicly available LLMs are English-centric.
 For instance, GPT-3’s training data consisted of approximately 93% English documents with the remaining 7% pertaining to other languages Brown et al.
@@
 One potential reason for this could be the “the curse of multilinguality” Conneau et al.
-(2020) which describes how as multilinguality increases, evaluation performance decreases.
-For example, multilingual LLMs have been shown to sacrifice performance on English tasks due to having to share a finite amount of pre-training tokens and model parameters across more languages Lin et al.
+(2020) which arises from having to share a finite model capacity across more languages Lin et al.
 (2022); Le Scao et al.
 (2022).
-Consequently, the cross-lingual transfer abilities of more performant, English-centric models is important for many NLP practitioners looking to deploy these models in real-world settings.
-A number of works have reported on the multilingual abilities of proprietary models such as GPT-3 and its derivatives across a range of NLU and NLG benchmarks Lai et al.
+Consequently, cross-lingual transfer abilities of performant English-centric models is highly valuable.
+Despite this, proprietary models such as GPT-3 and its derivatives have shown impressive performance in multilingual settings across a range of NLU and NLG benchmarks Lai et al.
 (2023a); Holmström et al.
 (2023); Armengol-Estapé et al.
-(2022).
-On tasks such as machine translation (MT), GPT-3.5-Turbo has been shown to outperform dedicated systems for some high-resource languages Hendy et al.
+(2022); Hendy et al.
 (2023); Lu et al.
@@
 (2023); Bang et al.
-(2023), particularly when translating between two non-English languages Laskar et al.
+(2023); Laskar et al.
 (2023).
-However, performance on low-resource languages generally lags behind.
-Meanwhile, other studies have focused on open-weight LLMs.
-Ye et al.
-(2023) compare multilingual reasoning capabilities of pretrained BLOOM Le Scao et al.
-(2023) and LLaMA Touvron et al.
-(2023a) and find that despite minimal amounts of non-English pretraining data, LLaMA has strong cross-lingual transfer abilities.
-Muennighoff et al.
-(2023) find that multilingual multitask finetuning improves BLOOM’s zero-shot abilities across all target languages compared with English-only finetuning.
-Finally, Chen et al.
-(2023) show that large-scale multilingual instruction tuning can improve performance on open-ended chat in multiple target languages.
-This last work is most closely related to our current work but differs in a few key factors.
-Firstly, we investigate multilingual chat capabilities of more recent English-centric models, pretrained with significantly more data.
-Secondly, inspired by Zhou et al.
-(2023) we adopt a less-is-more finetuning approach, controlling the amount and distribution of multilingual data.
-In doing so, we focus our analysis on the minimal amount of multilinguality needed to elicit cross-lingual transfer.
-Finally, our evaluations consider more target languages and distinct downstream tasks.
+Meanwhile, there is a growing body of research dedicated to studying similar capabilities of open-weight LLMs Bawden and Yvon (2023) and the benefits of multilingual tuning Ye et al.
+(2023); Muennighoff et al.
+(2023).
+For instance, large-scale multilingual instruction tuning has been shown to improve performance on open-ended chat tasks in multiple target languages for both English-centric and multilingual LLMs Chen et al.
+(2024); Li et al.
+(2023a); Weber et al.
+(2024).
+In concurrent work to ours, Shaham et al.
+(2024) report that this holds given minimal multilingual instruction tuning with the multilingual PaLM 2 model Anil et al.
+(2023).
+Our work closely relates to that of Chen et al.
+(2024) and Shaham et al.
+(2024) insofar as we explore how multilingual instruction tuning can improve cross-lingual transfer in LLMs.
+However, in contrast to these works, we focus on the minimal amount of multilinguality needed to elicit cross-lingual transfer in open-weight English-centric LLMs, and we evaluate on a range of tasks and target languages that cover the full distribution of a model’s pretraining data.
 3 Experimental Setup
-To explore the multilingual capabilities of English-centric LLMs, we instruction-tune a series of models on a fixed-size set of examples, varying the number of languages available.
-Following this, we evaluate the resulting models in multiple target languages on four distinct tasks that are representative of how LLMs may be used in downstream applications.
+To explore the multilingual capabilities of English-centric LLMs, we instruction tune a series of models on a fixed-size set of examples, varying the number of languages available.
+Following this, we evaluate the resulting models in multiple target languages on five distinct tasks that are representative of how LLMs may be used in downstream applications.
 3.1 English-centric LLMs
 A prevailing trend in the development of recent LLMs is a clear focus on scaling up the size of the pretraining corpus Hoffmann et al.
-(2022).
+(2022); Sardana and Frankle (2023).
 For instance, open-weight LLMs such as Falcon Almazrouei et al.
 (2023) and Llama 2 Touvron et al.
-(2023b) were trained on 1.5 and 2 trillion tokens respectively.
-Yet, although these numbers far surpass the 300 billion tokens used to train GPT-3 Brown et al.
-(2020), the distribution of language coverage remains similar across models with more than 90% pertaining to English (see Table 1).
-For our main experiments, we focus on Llama 2 7b, but also consider Llama 2 70b (§5.2) and Falcon 7b (Appendix E) to study the effect of model scaling and different training approaches respectively.
+(2023) were pretrained on 1.5 and 2 trillion tokens, respectively.
+Yet, while this far surpasses the 300 billion tokens used to train GPT-3 Brown et al.
+(2020), the distribution of language coverage remains similar with more than 90% pertaining to English (see Table 1).
+For our main experiments, we focus on Llama 2 7b.
+We then consider Llama 2 70b (§5.2) to study the effect of model scaling on multilinguality.
+Additionally, we also test whether our findings generalise to other LLMs using Falcon 7b and Llama 3 8b MetaAI (2024), which employ different pretraining approaches (see Appendices F and G).
 3.2 Instruction-tuning Data
@@
 Specifically, we sample 3,200 unique English instances as an initial monolingual dataset, which we refer to as ‘Mono’.
-To construct datasets for multilingual finetuning we sample 200 unique training examples from each of five most frequent non-English languages in Guanaco (Spanish, Russian, German, Chinese, French).
+To construct datasets for multilingual finetuning we sample 200 unique training examples from each of the five most frequent non-English languages in Guanaco (Spanish, Russian, German, Chinese, and French).
 Given these subsets, we incrementally substitute English examples in Mono for non-English ones, one language at a time, following the order of how frequently each language appears in Guanaco.
-The resulting multilingual datasets are denoted as Multi-$i$, where $i$ equals the number of distinct languages included.
+The resulting multilingual datasets are denoted as Multi-$i$, where $i$ indicates the number of distinct languages included.
+Table 3 provides a detailed summary of the makeup of these datasets.
 For comparison, we also train models on the full Guanaco dataset, which includes more than 30 distinct languages.
 3.3 Instruction-tuning and Inference Settings
-For efficient instruction tuning, we use LoRA adapters ($R=64$, $\alpha=16$), leveraging Hugging Face’s TRL library [^3] https://github.com/huggingface/trl.
-We train all models for 2,000 update steps, using an effective batch size of 64 and a learning rate of $1^{e-5}$.
-Sequences longer than 1024 tokens are truncated.
-For 7-billion parameter models, we use two NVIDIA GeForce RTX 3090 with 24GB of memory.
-The time required for each training run is approximately 8 hours.
-For the larger 70-billion parameter model, discussed in §5.2, we use the same hyperparameters and train on four NVIDIA A100 GPUs with 80GB of memory each.
-Here, a single training run takes approximately 20 hours.
-At inference time, we load the instruction-tuned models with vLLM Kwon et al.
+[…]
+For efficient inference, we employ vLLM Kwon et al.
 (2023).
-For open-ended generation tasks, we use top-p sampling (p=0.9) with a temperature of 0.8 and frequency penalty of 0.2.
-For the more constrained generation tasks we reduce the temperature to 0.001.
-To account for randomness under these sampling parameters, we report aggregated results for all tasks by repeating inference experiments with three random seeds.
+We use nucleus sampling Holtzman et al.
+(2019) ($p=0.9$) with a temperature of 0.8 for open-ended generation tasks and a temperature of 0.001 for the more constrained QA-style tasks to encourage more deterministic behaviour.
+For all tasks, we report the mean performance across three inference runs using different random seeds.
 3.4 Evaluation Tasks and Languages
-As evaluation tasks, we consider open-ended chat, extractive question answering, commonsense reasoning and natural language inference.
-As all of these tasks differ in terms of the availability and representation of ground-truth labels, we describe the specific evaluation strategies in the following section.
-Finally, we investigate performance on different target languages, which we select based on the makeup of Llama 2’s pretraining data (see Table 1).
-In doing so, we aim to investigate i) the supervised setting, where a target language has been seen during pretraining and finetuning; ii) the zero-shot cross-lingual setting, where a target language may have been seen during pretraining but not during finetuning Wu and Dredze (2019); and iii) the extreme low-resource setting, where the target language has not intentionally been seen at any stage of the training procedure.
-According to Touvron et al.
-(2023b), German (de), French (fr), Swedish (sv), Chinese (zh), Spanish (es) and Russian (ru) represent the most frequently seen non-English pretraining languages with an estimated 3.4 to 2.6 billion tokens.
-Catalan (ca), Norwegian (no) and Bulgarian (bg) represent languages less frequently seen during pretraining with an estimated 800 to 400 million tokens.
-And finally, we select Icelandic (is), Hindi (hi) and Greek (el) to represent extremely low-resource languages, whose frequency in Llama 2’s pretraining data is not available but may still appear in small amounts due to intentional or unintentional contamination Blevins and Zettlemoyer (2022).
+As evaluation tasks, we consider single-turn dialogue, sentence simplification, extractive question answering, commonsense reasoning and natural language inference.
+Since all of these tasks differ in terms of the availability and representation of ground-truth labels, we describe the specific evaluation strategies in the following section.
+Figure 2: Average helpfulness of single-turn dialogue responses from Llama 2 7b given incremental multilingual instruction tuning.
+Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+We select target languages based on the makeup of Llama 2’s pretraining data (see Table 1).
+In doing so, we aim to study the effect of multilingual instruction tuning in both supervised and 0-shot cross-lingual settings Wu and Dredze (2019), and report results on high-, medium-, and low-resource languages with regard to the pretrained LLM, covering distinct language families and scripts.
+German (de), French (fr), Swedish (sv), Chinese (zh), [^5] Note, Touvron et al.
+(2023) do not distinguish between Simplified and Traditional Chinese in their reporting.
+For our purposes, we explicitly use Simplified Chinese.
+Spanish (es), Russian (ru), and Italian (it) represent high-resource languages with an estimated 3.4 to 2.2 billion tokens.
+Portuguese (pt), Vietnamese (vi), Korean (ko), Finnish (fi), Hungarian (hu), Norwegian (no), [^6] Again, lacking specification between Norwegian Bokmål and Nynorsk in Touvron et al.
+(2023), we explicitly consider the former.
+Bulgarian (bg), and Slovenian (sl) represent medium-resource languages, with an estimated 1.8 to 0.2 billion tokens.
+And finally, we select Greek (el), Hindi (hi), and Icelandic (is) as low-resource languages, whose frequency in Llama 2’s pretraining data is not known but are likely appear in small amounts due to contamination Blevins and Zettlemoyer (2022).
 Unfortunately, existing benchmark evaluation datasets do not cover all of these languages and thus we limit the target languages in those tasks to the available subset of our target languages.
 4 Experiments
-4.1 Chat
+4.1 Single-turn Dialogue
 General-purpose chatbots are a popular application of instruction-tuned LLMs.
-To evaluate chat performance, we make use of the AlpacaEval prompt dataset Dubois et al.
+To evaluate performance in this type of setting, we make use of the AlpacaEval prompt dataset Dubois et al.
 (2023), which includes a diverse set of prompts for open-ended questions, creative writing, brainstorming, and other tasks.
@@
 As a translation engine, we follow Lai et al.
-(2023b) and use GPT-3.5-Turbo [^4] To translate chat prompts from AlpacaEval, we use gpt-3.5-turbo-0613..
-In contrast to using dedicated translation systems, employing GPT-3.5-Turbo for this purpose has the advantage of being able to explicitly specify instructions that allow for preserving code blocks, tables and terminology, which we include as part of our translation prompt (see Figure B).
+(2023b) and use GPT-3.5-Turbo.
+[^7] To translate prompts from AlpacaEval, we use gpt-3.5-turbo-0613.
+In contrast to using dedicated translation systems, employing GPT-3.5-Turbo for this purpose has the advantage of being able to explicitly specify instructions that allow for preserving code blocks, tables, and terminology, which we include as part of our translation prompt (see Figure B.1).
 Furthermore, since GPT-3.5-Turbo is trained on instruction- and conversational-style data, we expect it to perform well at translating in this domain.
-To automatically evaluate open-ended chat responses, we leverage LLM-as-a-judge (Zheng et al., 2023).
+To automatically evaluate open-ended responses, we leverage LLM-as-a-judge Zheng et al.
+(2023).
 Following Zhou et al.
-(2023), given an input prompt and model’s response, we ask GPT-3.5-Turbo [^5] For chat evaluation, we use gpt-3.5-turbo-1106 due to its longer context window and cheaper inference costs.
+(2023), given an input prompt and model’s response, we ask GPT-3.5-Turbo [^8] For evaluation, we use gpt-3.5-turbo-1106 due to its longer context window.
 to grade the helpfulness of the response on a 6-point Likert scale (see Figure 20 for the prompt used).
 For each evaluation instance, we provide the prompt and model-generated response directly in the target language, which we found to be on par with evaluating via first translating responses into English (cf.
-Hada et al., 2023) (see Appendix C for more details).
+Hada et al., 2024, see Appendix E for more details).
 As noted by Chen et al.
 (2024), GPT-3.5-Turbo sometimes ignores the fact that the output language differs from the input language.
-Therefore, we force a score of 1 (indicating least helpful) if the language of the response does not match the intended target language.
+Therefore, we force a score of 1 (indicating least helpful) if the language of the response does not match the intended target language according to OpenLID Burchell et al.
+(2023).
 Results
-Figure 2: Average helpfulness of chat responses from Llama 2 7b with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
-Figure 3: X-QuAD results for Llama 2 7b with incremental multilingual instruction tuning.
-Results are shown for both multilingual prompting (en:xx) and monolingual prompting (xx:xx).
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
-Figure 2 shows the helpfulness scores assigned by GPT-3.5-Turbo for all models for each of the target languages.
-For English, we observe a reasonably strong and stable performance across all incremental multilingual instruction tuning settings, indicating that chat performance does not diminish with the increase in non-English training examples.
-In contrast, for most non-English target languages, performance increases significantly when moving from monolingual to bilingual instruction tuning.
-Most notably though, performance tends to plateau when instruction tuning with as few as three different languages.
-This observation holds for languages in both the supervised setting and the zero-shot setting, indicating that three finetuning languages is sufficient to elicit zero-shot cross-lingual transfer on this task.
-For languages in the extremely low-resource category, performance remains low despite multilingual finetuning.
-Manual inspection of these outputs reveal that, while these responses generally look convincing at first glance and are sufficient for language identification (see Figure 1), they are mostly nonsensical.
-This suggests that useful zero-shot cross-lingual transfer abilities of Llama 2 7b are limited to languages seen with higher frequency during pretraining.
+Figure 3: SARI weighted by IO language agreement for Llama 2 7b given incremental multilingual instruction tuning.
+Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx).
+Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 2 shows the helpfulness scores assigned by the LLM judge on 16 target languages.
+For English, performance remains uniform across all multilingual instruction tuning settings.
+For high- and medium-resource non-English target languages, performance increases significantly when moving from monolingual to bilingual instruction tuning (Multi-2).
+Meanwhile, performance plateaus when training with as few as three languages (Multi-3), with no substantial differences observed between supervised and 0-shot settings.
+For low-resource languages, performance remains low despite multilingual finetuning.
+Manual inspection reveals these outputs, while initially convincing and sufficient for language identification (see Figure 1), are mostly nonsensical.
+These results indicate that instruction tuning with as few as two to three languages is necessary and sufficient to elicit 0-shot cross-lingual generalisation among high- and medium-resource languages on this task, with language exposure during pretraining being the main limitation to this generalisation ability.
+4.2 Sentence Simplification
+Sentence simplification aims to make complex sentences easier to read and understand.
+Due to a lack of high-quality supervised training data, automatic sentence simplification remains a challenging task Štajner (2021), and thus stands to benefit from the few- and 0-shot generalisation capabilities of instruction-following LLMs.
+To assess performance on this task, we use the MultiSim benchmark Ryan et al.
+(2023), which includes sentence-aligned parallel datasets in multiple languages.
+Since the individual datasets in MultiSim are taken from distinct sources, the amount of available data varies across languages.
+We sample 1,371 complex-simple sentence pairs for each language, except for Slovenian, where we use all 939 instances available (see Appendix D for details).
+To prompt models for sentence simplification, we repurpose the detailed instructions given to crowdworkers for the creation of the ASSET corpus Alva-Manchego et al.
+(2020), which has been shown to be effective in few-shot settings Kew et al.
+(2023).
+In its original form, this prompt explicitly states that the output should be suitable for “learners of English”.
+We remove this language specification so that the model is not explicitly instructed on the target language and instead is expected to infer it from the complex input sentence provided.
+To translate the prompt into each of our considered target languages, we used a free-tier machine translation service.
+[^9] https://www.deepl.com/translator We also consider both a cross-lingual prompt setting (en:xx) and a monolingual prompt setting (xx:xx).
+The former presents the task instruction in English and provides the complex source sentence in the relevant target language, while the latter is entirely in the target language.
+An example of the prompt template is shown in Table 4.
+For evaluation, we report SARI Xu et al.
+(2016) as an indicator of simplicity and adequacy given the reference simplification.
+To account for IO language agreement, we weight a model’s corpus-level SARI scores by the percentage of IO language agreement it achieved.
+This penalises model if it consistently produces outputs in a language that does not match the source sentence.
+For example, if the model generates English outputs given German input sentences at a rate of 50%, the final score is reduced by half.
+Figure 4: XQuAD results for Llama 2 7b given incremental multilingual instruction tuning.
+Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx).
+Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Results
+Figure 3 shows the performance according to IO-weighted SARI on each of the target languages considered given incremental multilingual instruction tuning.
+Similar to the results on the single-turn dialogue task, performance remains uniform for English as the ratio of multilingual instructions increases.
+For all non-English target languages, we observe a large discrepancy between the cross-lingual prompting strategy and the monolingual prompting strategy.
+While the former tends to improve as the number of languages increases, it fails to match that of the monolingual prompting strategy under all conditions.
+In contrast, the results for the monolingual prompting strategy closely resemble those of the single-turn dialogue task across all available target languages: performance gains are most pronounced when moving from monolingual to bilingual instruction tuning, and they generally plateau with as few as three instruction tuning languages.
 4.3 Extractive Question Answering
-Figure 4: X-CSQA results for Llama 2 7b with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
-Figure 5: XNLI results for Llama 2 7b with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
 In contrast to open-ended questions commonly used to query LLMs in chat settings, extractive question answering requires the model to identify relevant answer spans within longer context passages provided as part of the prompt.
-This kind of task closely resembles a retrieval augmented generation (RAG) setting, which is a popular method for extending an LLM’s knowledge with additional data not available during training Lewis et al.
+Such a task closely resembles a retrieval augmented generation (RAG) setting, which is a popular method for extending an LLM’s knowledge with additional data not available during training Lewis et al.
 (2020); Izacard and Grave (2021).
-To evaluate our models on this task in multiple target languages, we use XQuAD Artetxe et al.
+To evaluate our models on this task in multiple target languages we use XQuAD Artetxe et al.
 (2020).
@@
 This provides 1,190 QA pairs that were professionally translated into different languages.
-For this task, we consider both a multilingual prompt setting and a monolingual prompt setting.
-The former presents the task instruction in English and provides the context passage and question in the relevant target language (en:xx), while the latter is entirely in the target language (xx:xx).
 As a starting point, we borrow the English prompt from Lai et al.
@@
 This allows us to better isolate the relevant answer string in the generative model’s output.
-A response is considered correct if the model’s answer string matches the ground truth answer after minimal postprocessing.
-[^7] We find that some postprocessing of model outputs is necessary for certain languages.
-Specifically, when queried with German and Russian prompts, all models consistently repeated the question before providing the extracted answer.
-To handle such edge cases, we strip away the question and truncate the system output to a maximum of 50 characters or the first line break, whichever comes first.
+A response is considered correct if the ground truth answer can be matched with the beginning of the model’s response after minimal post-processing.
+[^11] We find that some post-processing of model outputs is necessary for certain languages.
+Specifically, when queried with German and Russian prompts, models consistently repeated the question before providing the extracted answer.
+To handle such cases, we strip away the repeated question and truncate the response to a maximum of 50 characters or the first line break, whichever comes first.
+Again, we consider both cross-lingual (en:xx) and monolingual (xx:xx) prompting strategies.
 An example of our prompting strategy for this task and model outputs is shown in Table 5.
@@
 Figure 4 shows the performance on each of our target languages in XQuAD given incremental multilingual instruction tuning.
-Similar to the results on the chat task (§4.1), we observe no performance degradation on English as the ratio of multilingual instructions increases.
-For most non-English target languages in the supervised setting, we see that when using the multilingual prompting strategy (en:xx), performance remains relatively uniform, with the exception of Chinese.
-In contrast, given the monolingual prompting strategy (xx:xx), multilingual instruction tuning can lead to small but noticeable gains in performance over monolingual tuning.
-Specifically, for German and Chinese, performance improves when moving from monolingual to bilingual instruction tuning and again plateaus rather quickly with as few as three languages.
-For languages in the extremely low-resource setting (Hindi and Greek), multilingual instruction tuning fails to improve the model’s ability to handle these languages.
-While performance gains on this task are generally less pronounced than on the open-ended chat task, we note that zero-shot extractive QA is inherently challenging for LLMs tuned on conversational instructions as they tend to generate verbose responses rather than the single word or entity that correctly matches the ground truth.
+Again, we observe consistent performance for English under multilingual instruction tuning settings.
+Meanwhile, performance on high-resource languages such as German and Chinese exhibits marginal gains when moving from monolingual to bilingual instruction tuning and plateaus with relatively few languages.
+For low-resource languages (e.g., Hindi and Greek), performance remains consistently low across all settings, indicating that the model’s ability to generalise to these languages is limited by the lack of exposure during pretraining.
+Comparing the cross-lingual and monolingual prompting strategies, we observe that the latter generally leads to better performance.
+This agrees with the results from the sentence simplification task (§4.2) and highlights the benefit of prompting under an IO language agreement assumption.
+While performance gains on this task are generally less pronounced than on the previous tasks considered, we note that 0-shot extractive QA is inherently challenging for LLMs tuned on conversational instructions as they tend to generate verbose responses rather than the single word or entity that correctly matches the ground truth.
 4.4 Commonsense Reasoning
-Effectively understanding natural language requires some representation of the natural world and how concepts can relate to one another.
-Commonsense knowledge describes the set of general facts that reflects this.
-To evaluate how well English-centric models can reason across languages, we use the X-CSQA dataset.
+Figure 5: X-CSQA results for Llama 2 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 6: XNLI results for Llama 2 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+‘Commonsense knowledge’ is the term frequently used to refer to the set of general facts that reflect how concepts can relate to one another in the real world Li et al.
+(2022); Talmor et al.
+(2019).
+Effectively understanding natural language requires some representation of these concepts and relationships, making commonsense reasoning a key skill for LLMs.
+To evaluate how well English-centric models can apply commonsense reasoning across languages, we use the X-CSQA dataset Lin et al.
+(2021).
 [^12] We report results for X-CSQA measured on the validation set of 1,000 questions.
-This dataset consists of questions paired with multiple choice answers which aim at assessing general, language-agnostic world knowledge involving different types of commonsense reasoning Talmor et al.
-(2019); Lin et al.
-(2021).
-Given a question and a set of five possible answers from A-E, we prompt the model to output the letter corresponding to the answer that best answers the question.
+This dataset contains questions paired with multiple choice answers which aim at assessing general, language-agnostic world knowledge involving different types of commonsense reasoning.
+Given a question and a set of five possible answers from A-E, we prompt the model to output the letter corresponding to the most suitable answer.
 Again, we borrow the English prompt template for this task from Lai et al.
 (2023a) as a starting point and translate it into each target language.
+For this task we consider only the monolingual prompting strategy (xx:xx), which proved most effective in the previous tasks.
 An example of the prompt used is given in Table 6.
-Finally we assess performance with the monolingual prompting strategy (xx:xx).
 Results
@@
 Again, for English, we observe that multilingual instruction tuning does not significantly degrade performance.
-However, for non-English target languages, incremental multilingual instruction tuning fails to deliver improvements over monolingual tuning.
-This contrasts with the results on the previous tasks considered.
-Figure 6: Chat performance of Llama 2 7b given multilingual instruction tuning using native non-English (Native NE) examples compared to using translated non-English (MT NE) examples.
-Error bars show a confidence interval of 95%.
+However, in contrast to the previous tasks considered, we do not observe any consistent gains for non-English target languages on the basis of multilingual instruction tuning, suggesting that there are limitations on what tasks this benefits.
+Figure 7: Comparison of Llama 2 7b given multilingual instruction tuning using native non-English (Native NE) examples vs.
+translated non-English (MT NE) examples.
+Error bars show a 95% confidence interval.
 4.5 XNLI
-Natural language inference (NLI) is an important skill for LLMs, especially as input and output sequences grow to consist of multiple sentences.
-Given two sentences, this task aims to recognise a relationship between them either as entailment, contradiction or neutral.
-To assess an LLM’s ability to solve this task given multilingual instruction tuning, we use XNLI (Conneau et al., 2018) and evaluate performance on the official test split.
-For this task, we use the implementation in the lm-evaluation harness Gao et al.
-(2023) [^9] https://github.com/EleutherAI/lm-evaluation-harness, which uses language-specific prompts for each target language (i.e.
-monolingual prompting (xx:xx)).
-Instead of querying the model to generate the desired label, multiple queries are constructed for each test instance (one for each possible label) and scored by the model.
-The sequence with the highest likelihood is chosen as the model’s answer.
+Natural language inference (NLI) is an important skill for LMs, as it ensures textual coherence, which is particularly critical as input and output sequences grow in length.
+Given two sentences, this task aims to recognise a relationship between them either as entailment, contradiction, or neutral.
+To assess an LLM’s ability to solve this task given multilingual instruction tuning, we use XNLI Conneau et al.
+(2018) and evaluate performance on the official test split.
+For this task, we use the implementation in the LM Evaluation Harness Gao et al.
+(2023).
+[^13] https://github.com/EleutherAI/lm-evaluation-harness Instead of querying the model to generate the desired label, we use rank classification Sanh et al.
+(2022); Brown et al.
+(2020), which allows for direct comparison to the base model.
+In this setting, multiple queries are constructed for each test instance given a predefined template and subsequently scored by the model.
+The most probable query sequence under the model is then chosen as the final answer.
+Prompt templates for this task are language-specific making it equivalent to monolingual prompting (xx:xx).
+An example of this task is provided in Table 7.
 Results
-Figure 5 shows the accuracy measured on our target languages from XNLI given incremental multilingual instruction tuning.
-Strikingly, performance for all target languages remains uniform regardless of the amount of multilinguality used in finetuning.
-Furthermore, the equal performance between the Guanaco baseline and our Mono and Multi-$i$ finetuned models suggests that scaling up instruction diversity and the number of languages seen during finetuning beyond just six languages is still insufficient for delivering noticeable gains on this task.
+Figure 6 shows the accuracy measured on our target languages from XNLI.
+Strikingly, performance for all target languages remains uniform regardless of the amount of multilinguality used during finetuning.
+Furthermore, the comparable performance across all training settings suggests that, in general, instruction tuning fails to provide measurable gains in this challenging task.
 5 Further Analysis
-The results on four distinct evaluation tasks show that multilingual instruction tuning benefits open-ended chat in non-English languages most strongly (§4.1).
-In this section, we focus on this task to investigate the impact of instruction diversity and model scaling.
+Our experimental results on five distinct tasks demonstrate that multilingual instruction tuning benefits open-ended single-turn dialogue and sentence simplification in non-English languages most strongly.
+In this section, we focus on single-turn dialogue to investigate the impact of instruction diversity vs.
+language diversity and model scaling.
 5.1 Instruction Diversity vs. Language Diversity
-Figure 7: Chat performance of Llama 2 70b (denoted by triangular points) and Llama 2 7b (semi-transparent bars) with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
+Figure 8: Average helpfulness of single-turn dialogue responses from Llama 2 70b (denoted by triangular points) and Llama 2 7b (semi-transparent bars) given incremental multilingual instruction tuning.
+Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
 Diversity is a key factor for sample efficient instruction tuning Zhou et al.
@@
 Since our experiments in §4 make use of native non-English training instances from the OpenAssistant dataset, a potential confounding factor could be that adding more languages during finetuning also introduces more diverse training instructions.
-To investigate this, we retrain Multi-$i$ models using translated instruction-tuning data from Mono in place of native non-English examples, following the same incremental recipe as described in §3.2.
+To investigate this, we retrain Multi-$i$ models using translated instruction tuning data from Mono in place of native non-English examples, following the same incremental recipe as described in §3.2.
 This ensures that the data distribution remains constant as multilinguality increases.
-As a translation engine, we again use GPT-3.5-Turbo [^10] Since conversational training instances can be quite long, sometimes exceeding the default 4k token context window, we use gpt-3.5-tubo-16k for this translation task.
+As a translation engine, we again use GPT-3.5-Turbo141414 Since conversational training instances can be quite long, sometimes exceeding the default 4k token context window, we use gpt-3.5-tubo-1106 for this translation task due to its longer context window (16k).
 and the prompt template provided in Figure 10.
-Figure 6 compares the resulting performance on the open-ended chat task on a subset of our target languages (results for the remaining target languages are provided in Figure 15).
-Notably, we observe no significant differences between tuning with distinctly native non-English examples compared to those derived via automatic translation.
-This indicates that the gains attributed to increased multilinguality are not conflated with an increase in the diversity of instructions.
+Figure 7 compares the resulting performance on the single-turn dialogue task for a subset of our target languages (results for the remaining target languages are provided in Figure 19).
+Notably, we observe no significant differences between tuning with distinctly native non-English examples compared to those derived via automatic translation, indicating that gains attributed to increased multilinguality are not conflated with an increase in the diversity of instructions.
+Following this, in Appendix H, we dive deeper to investigate the role of language diversity given a fixed budget of non-English training examples.
+From these ablation experiments, we find that the impact of increasing the number of languages is stronger than that of increasing the number of examples for a given language.
 5.2 Scaling up Model Size
-In order to investigate the effect of model scaling, we also repeat our chat experiments using Llama 2 70b as the underlying base model.
-Figure 7 shows the resulting helpfulness scores assigned by the LLM judge.
-Most notably, performance on non-English languages seen relatively frequently during pretraining is dramatically improved, often matching that of English.
-Secondly, we observe that the larger model’s performance on most non-English target languages tends to plateau with just two finetuning languages, unlike the smaller model that required three.
-Finally, while the performance on languages in the extremely low-resource scenario remains low, it exhibits a substantial relative improvement compared to the 7-billion parameter model.
-These results indicate that model scaling is extremely beneficial for exploiting the multilingual capabilities in English-centric models.
-This aligns with the assertion of Shaham et al.
-(2023) that larger models are more adept at handling multilinguality.
+In order to investigate the effect of model scaling, we repeat our single-turn dialogue experiments using Llama 2 70b as the underlying base model.
+Figure 8 shows the resulting helpfulness scores assigned by the LLM judge, with a direct comparison to Llama 2 7b.
+Most notably, performance on high- and medium-resource non-English target languages is dramatically improved, often matching that of English.
+However, despite these gains, multilingual instruction tuning with two languages remains essential to elicit cross-lingual generalisation on this task, further underpinning our main findings.
+In addition, we observe that the larger model’s performance on most non-English target languages tends to plateau with just two finetuning languages, unlike the smaller model that typically required three.
+Finally, while the performance on low-resource languages remains poor, it exhibits a substantial relative improvement compared to the 7-billion-parameter model.
+These results indicate that model scaling is extremely beneficial for exploiting the multilingual capabilities in English-centric models and aligns with findings from Shaham et al.
+(2023).
 6 Discussion
-Our findings show that multilingual instruction tuning can elicit zero-shot cross-lingual transfer in English-centric models, though its effectiveness on downstream performance varies across tasks.
-Notably, we observe significant performance gains for open-ended chat and noticeable gains on extractive QA with monolingual prompting.
-In comparison, for highly structured tasks such as X-CSQA and XNLI that impose a strict constraint on the output space regardless of the input language (e.g., choosing from options ‘A’, ‘B’, ‘C’, etc.), multilingual instruction tuning does not seem to provide substantial benefits.
-This distinction highlights that multilingual instruction tuning is most beneficial for open-ended generative tasks where there is an implicit agreement between the input query language and the language output by the model.
-We posit that our findings align with the superficial alignment hypothesis Zhou et al.
-(2023).
-While instruction tuning guides the model towards a desirable ‘subdistribution of formats’ to use, a small amount of multilingual instruction data allows the model to learn a simple mapping between input and output languages.
-In essence, it helps steer the model towards language-specific ‘subdistributions of formats’ that agree with the language of the input.
-Therefore, we conclude that multilingual instruction tuning is of most benefit in scenarios where IO language agreement is required.
-Most surprisingly, we find that multilingual instruction tuning with just three languages is sufficient to encourage IO language agreement in smaller LLMs, resulting in improved cross-lingual transfer on generative tasks.
-As model size increases, however, the number languages required decreases from three to two, indicating that larger models can learn this mapping easier.
-In both model sizes, adding more instruction tuning languages – including the target language itself – provides no significant gains.
-Despite benefits from cross-lingual transfer, our results highlight a clear gap in performance between English and non-English target languages on all tasks.
-This indicates that small-scale multilingual instruction tuning is insufficient to improve the intrinsic multilingual capabilities of English-centric LLMs and that more advanced interventions may be required (e.g.
-Pfeiffer et al., 2020; ImaniGooghari et al., 2023).
+Our findings show that multilingual instruction tuning aids cross-lingual transfer in English-centric LLMs, though its effectiveness on downstream performance varies across tasks and the degree to which a specific target language is seen during pretraining.
+For high- and medium-resource languages, we observe substantial performance gains on single-turn dialogue and sentence simplification with monolingual prompting, as well as marginal gains on extractive QA with monolingual prompting.
+Meanwhile, for low-resource languages and highly structured tasks that impose a strict constraint on the output space regardless of the input language (e.g., multiple choice QA), multilingual instruction tuning has little impact.
+This distinction highlights that multilingual instruction tuning is most beneficial for more open-ended generative tasks that assume IO language agreement, and crucially, that this benefit is limited to high- and medium-resource languages with respect to model’s pretraining language distribution.
+Most surprisingly, we find that multilingual instruction tuning with just two to three languages is necessary and sufficient to promote this generalisation in all four of the English-centric LLMs considered.
+Furthermore, we consistently see that performance tends to plateau rather quickly and adding more instruction tuning languages – including the target language itself – typically provides no significant gains.
+This underscores that the performance gains are largely the result of cross-lingual transfer and suggests that there may be diminishing returns associated with scaling up multilingual instruction tuning beyond just a handful of languages.
+We posit that our findings align with the superficial alignment hypothesis put forward by Zhou et al.
+(2023), which states that a model acquires its knowledge and abilities during pretraining, while instruction tuning simply guides the model towards a desirable ‘subdistribution of formats’ to use when prompted.
+We refine this hypothesis by adding that just a small amount of multilingual instruction data also encourages the model to learn a simple mapping between input and output language.
+This guides the model towards language-specific subdistributions, leading to better performance on tasks where IO language agreement is required.
 7 Conclusion
-In this paper, we investigated the use of multilingual instruction tuning to elicit multilingual capabilities of English-centric LLMs.
-Our results show that finetuning with as few as three languages can promote cross-lingual transfer and allow models to better exploit the relatively small amounts of non-English data seen during pretraining.
-Experiments conducted on four distinct tasks revealed that this can lead to significant performance improvements on open-ended generative tasks that assume input and output language agreement.
-Furthermore, we observed that these effects are amplified in larger models, requiring only two finetuning languages for successful multilingual instruction tuning.
-The effectiveness of cross-lingual transfer, which reduces the need for extensive multilingual instruction tuning, is good news for LLM developers.
-However, future work could explore other methods to encourage smaller models to better leverage the limited non-English data seen during pretraining and whether there are tasks for which language-specific instruction tuning is of higher importance.
-8 Limitations
-This work focused on detecting and evaluating non-English responses from English-centric LLMs.
-Our experiments demonstrated that cross-lingual transfer, which can be elicited given small amount of multilingual instruction tuning, can be effective for improvinging the non-English performance of English-centric models.
-That said, our analysis focused primarily on Indo-European languages, except for Chinese.
-As a consequence, it remains unclear how well our results generalise to more distant or topologically diverse languages, however, performance gains observed on Chinese suggest that generalisation is more dependent on exposure during pretraining than language similarity.
-For our experiments and analysis on model training data, we relied on automatic language identification.
+We investigated the minimal amount of multilingual instruction tuning required to elicit multilingual capabilities of four distinct English-centric LLMs.
+Our results show that finetuning with as few as two to three languages is necessary and sufficient to promote effective cross-lingual transfer, allowing models to better exploit the relatively small amounts of non-English data seen during pretraining.
+Experiments on five distinct tasks revealed that this can lead to significant performance improvements for high- and medium-resource languages on open-ended generative tasks that assume input/output language agreement.
+While the effectiveness of cross-lingual transfer is indeed good news for LLM developers, future work could explore methods to reduce the performance gap between English and non-English languages, particularly for smaller models, as well as investigating tasks for which language-specific instruction tuning may be of greater importance, such as tasks involving cultural awareness.
+Limitations
+Our experimental results demonstrate that a relatively small amount of multilingual instruction tuning data can elicit highly valuable cross-lingual transfer, leading to improved performance in open-ended generation tasks that assume IO language agreement.
+However, we have not explored tasks involving more complicated mappings between languages, such as tasks involving extensive cross-lingual processing or translation.
+In this work we rely on automatic language identification to construct multilingual training data and in evaluating model outputs.
 To this end, we employed the OpenLID model from Burchell et al.
@@
 Despite low error rates achieved by this model, language identification is not perfect and can lead to some texts being misidentified.
-To mitigate the risk of language contamination Blevins and Zettlemoyer (2022) in our controlled experiments, we only included training examples whose language is identified with a confidence threshold $\geq$ 0.8.
-In §5.1 we investigated the impact of multilingual diversity versus training example diversity.
-While our findings revealed that there is no significant difference between these two settings, we note that even when finetuning with the original native non-English examples, task diversity may be inherently limited by design of the data collection.
+To mitigate the risk of unintentional language contamination Blevins and Zettlemoyer (2022) in our finetuning datasets we include training examples whose language is identified with a confidence threshold $\geq$ 0.8.
+In §5.1 we investigate the impact of multilingual diversity versus training example diversity.
+While our findings reveal that there is no significant difference between these two settings, we note that even when finetuning with the original native non-English examples, task diversity may be inherently limited by design of the data collection.
 For instance, regardless of the language used, crowdworkers were asked to follow the same set of guidelines [^15] https://projects.laion.ai/Open-Assistant/docs/guides/guidelines when creating the data.
+Finally, to evaluate the quality of model-generated responses in the single-turn dialogue task, we rely on LLM-based evaluation methods.
+While a number of works have shown that LLM-based evaluation provides a decent proxy for assessing the quality of dialogue responses, achieving strong correlations with human judgements Liu et al.
+(2023); Kocmi and Federmann (2023); Zheng et al.
+(2023), we stop short of empirically establishing this agreement on our own model outputs.
+To ensure the validity of model outputs we manually assessed the adequacy of randomly sampled responses in three target languages (en, es, and de).
+This assessment revealed a similar trend to our LLM-based evaluation, namely that Spanish-language outputs were generally slightly worse than their English-language counterparts but slightly better than those in German.
+Ethical Considerations and Risks
+This work aims to evaluate multilingual capabilities of English-centric LLMs.
+In doing so, we acknowledge potential ethical considerations and risks associated with our research.
+Firstly, LLMs have been shown to inadvertently perpetuate biases present in their training data, which can lead to unexpected and unfair outcomes when these models are used in real-world applications.
+Therefore, measures must be taken to minimise this risk (e.g., using additional alignment strategies, rigorous testing in multilingual settings) before deploying public-facing models.
+Secondly, when building on top of English-centric LLMs, there is a risk of cultural homogenization, where nuances and diversity of different languages could potentially be lost due to their under-representation in the training data.
 Acknowledgements
-We sincerely thank our friends and colleagues, Fabian Aiolfi, Thea Bächler, Anastassia Shaitarova, Finnur Ágúst Ingimundarson, Cui Ding, Andrianos Michail, Janine Aeberhard, Arnisa Fazla, Farhad Nooralahzadeh, Omnia Ibrahim, Xuan Lam and Manasi Muglikar, for helping us with language-specific questions and validating translations used in our experiments, as well as Lena Bolliger and Patrick Haller for providing valuable feedback.
+We sincerely thank our friends and colleagues, Fabian Aiolfi, Thea Bächler, Anastassia Shaitarova, Finnur Ágúst Ingimundarson, Cui Ding, Andrianos Michail, Janine Aeberhard, Arnisa Fazla, Farhad Nooralahzadeh, Omnia Ibrahim, Xuan Lam, Manasi Muglikar, Anne Göhring, and Alessia Battisti for helping us with language-specific questions and validating translations used in our experiments, as well as Lena Sophia Bolliger and Patrick Haller for providing valuable feedback on earlier versions of this paper.
 Rico Sennrich acknowledges funding by the Swiss National Science Foundation (project MUTAMUR; no.
@@
 Appendix A Pretraining Data for English-centric LLMs
-LLMs are pretrained on enormous amounts of unlabelled text gathered from existing corpora and the web.
-With the notable exception of BLOOM (Le Scao et al., 2023), very limited information is offered about the distribution of languages represented in datasets used to train LLMs and the subsequent performance on non-English languages.
+With the notable exception of BLOOM Le Scao et al.
+(2023) and the recent OLMo models Groeneveld et al.
+(2024), very limited information is offered about the distribution of languages represented in datasets used to train LLMs and the subsequent performance on non-English languages.
 Table 1 provides an overview of the document-level language distributions of the LLMs used in this paper.
 For Llama 2, the information is taken from the original paper Touvron et al.
-(2023b), in which the authors analyse the training data using a fastText language classifier on corpus documents with a threshold of 0.5.
-For GPT-3, we use the official dataset statistics made available on Github [^12] https://github.com/openai/gpt-3/blob/master/dataset_statistics/languages_by_document_count.csv which provide document-level language identification information.
-To gather statistics for Falcon, we inspected a sample of the RefinedWeb dataset Penedo et al.
-(2023a) that was constructed to train this family of models.
+(2023), in which the authors analyse the training data using a fastText language classifier on corpus documents with a threshold of 0.5.
+For GPT-3, we use the official dataset statistics made available on GitHub which provide document-level language identification information.
+[^16] https://github.com/openai/gpt-3/blob/master/dataset_statistics/languages_by_document_count.csv To gather statistics for Falcon, we inspected a sample of the RefinedWeb dataset Penedo et al.
+(2023) that was constructed to train this family of models.
 [^17] https://huggingface.co/datasets/tiiuae/falcon-refinedweb Using the OpenLID fastText model from Burchell et al.
 (2023), we identify the most frequent document-level languages based on approximately 320 million examples from this corpus.
-Languages are identified with a confidence threshold of $\geq$ 0.6 and predictions below this threshold are aggregated under ‘unknown’.
+Following Touvron et al.
+(2023), languages are identified based on a confidence threshold of 0.5 and predictions below this threshold are aggregated under ‘unknown’.
+Note, that while this language distribution has a higher concentration of English data than both Llama 2 and GPT-3, Almazrouei et al.
+(2023) combine the RefinedWeb with additional curated corpora to train the Falcon models, and thus the true language distributions for this model may differ from these estimates.
 Llama 2 Falcon GPT-3
 # Language Percent # Language Percent # Language Percent
-1 English 89.70% 1 English 94.790% 1 English 93.69%
-2 unknown 8.38% 2 unknown 4.944% 2 German 1.20%
-3 German 0.17% 3 Standard Malay 0.031% 3 French 1.02%
-4 French 0.16% 4 Swahili 0.027% 4 Portuguese 0.64%
-5 Swedish 0.15% 5 French 0.016% 5 Italian 0.58%
-6 Chinese 0.13% 6 Portuguese 0.012% 6 Spanish 0.51%
-7 Spanish 0.13% 7 Chinese (Traditional) 0.011% 7 Dutch 0.37%
-8 Russian 0.13% 8 Sardinian 0.011% 8 Polish 0.25%
-9 Dutch 0.12% 9 Dutch 0.009% 9 Japanese 0.25%
-10 Italian 0.11% 10 Spanish 0.009% 10 Danish 0.16%
-11 Japanese 0.10% 11 Italian 0.008% 11 Norwegian 0.15%
-12 Polish 0.09% 12 German 0.008% 12 Romanian 0.13%
-13 Portuguese 0.09% 13 Danish 0.005% 13 Finnish 0.13%
-14 Vietnamese 0.08% 14 Russian 0.005% 14 Chinese (Simplified) 0.12%
-15 Ukranian 0.07% 15 Indonesian 0.005% 15 Russian 0.11%
-16 Korean 0.06% 16 Swedish 0.005% 16 Czech 0.10%
-17 Catalan 0.04% 17 Somali 0.005% 17 Swedish 0.06%
-18 Serbian 0.04% 18 Vietnamese 0.004% 18 Hungarian 0.06%
-19 Bahasa Indonesian 0.03% 19 Venetian 0.004% 19 Chinese (Traditional) 0.04%
-20 Czech 0.03% 20 Chinese (Simplified) 0.004% 20 Bahasa Indonesian 0.04%
-21 Finnish 0.03% 21 Tagalog 0.004% 21 Croation 0.04%
-22 Hungarian 0.03% 22 Northern Uzbek 0.003% 22 Turkish 0.04%
-23 Norwegian 0.03% 23 Limburgish 0.003% 23 Catalan 0.03%
-24 Romanian 0.03% 24 Polish 0.003% 24 Vietnamese 0.03%
-25 Bulgarian 0.02% 25 Turkish 0.003% 25 Slovenian 0.03%
-26 Danish 0.02% 26 Norwegian Nynorsk 0.003% 26 Estonian 0.02%
-27 Slovenian 0.01% 27 Norwegian Bokmal 0.003% 27 Slovak 0.02%
-28 Croatian 0.01% 28 Romanian 0.002% 28 Korean 0.02%
-n/a Icelandic n/a 30 Greek 0.002% 31 Greek 0.017%
-n/a Hindi n/a 60 Bulgarian 0.001% 45 Hindi 0.004%
-n/a Greek n/a 72 Icelandic 0.001% 47 Icelandic 0.004%
+1 English 89.70% 1 English 95.751% 1 English 93.69%
+2 unknown 8.38% 2 unknown 3.902% 2 German 1.20%
+3 German 0.17% 3 Standard Malay 0.045% 3 French 1.02%
+4 French 0.16% 4 Swahili 0.038% 4 Portuguese 0.64%
+5 Swedish 0.15% 5 French 0.019% 5 Italian 0.58%
+6 Chinese 0.13% 6 Sardinian 0.014% 6 Spanish 0.51%
+7 Spanish 0.13% 7 Portuguese 0.014% 7 Dutch 0.37%
+8 Russian 0.13% 8 Dutch 0.01% 8 Polish 0.25%
+9 Dutch 0.12% 9 Chinese (Traditional) 0.013% 9 Japanese 0.25%
+10 Italian 0.11% 10 Spanish 0.010% 10 Danish 0.16%
+11 Japanese 0.10% 11 Italian 0.010% 11 Norwegian 0.15%
+12 Polish 0.09% 12 German 0.009% 12 Romanian 0.13%
+13 Portuguese 0.09% 13 Danish 0.008% 13 Finnish 0.13%
+14 Vietnamese 0.08% 14 Indonesian 0.007% 14 Chinese (Simplified) 0.12%
+15 Ukranian 0.07% 15 Somali 0.006% 15 Russian 0.11%
+16 Korean 0.06% 16 Swedish 0.006% 16 Czech 0.10%
+17 Catalan 0.04% 17 Russian 0.006% 17 Swedish 0.06%
+18 Serbian 0.04% 18 Venetian 0.005% 18 Hungarian 0.06%
+19 Bahasa Indonesian 0.03% 19 Vietnamese 0.005% 19 Chinese (Traditional) 0.04%
+20 Czech 0.03% 20 Northern Uzbek 0.005% 20 Bahasa Indonesian 0.04%
+21 Finnish 0.03% 21 Limburgish 0.005% 21 Croation 0.04%
+22 Hungarian 0.03% 22 Tagalog 0.005% 22 Turkish 0.04%
+23 Norwegian 0.03% 23 Chinese (Simplified) 0.005% 23 Catalan 0.03%
+24 Romanian 0.03%
+24 Vietnamese 0.03%
+25 Bulgarian 0.02% 27 Norwegian Bokmal 0.003% 25 Slovenian 0.03%
+26 Danish 0.02%
+26 Estonian 0.02%
+27 Slovenian 0.01% 33 Finnish 0.003% 27 Slovak 0.02%
+28 Croatian 0.01%
+28 Korean 0.02%
+35 Greek 0.002%
+n/a Greek n/a
+31 Greek 0.017%
+46 Hungarian 0.002%
+n/a Hindi n/a
+45 Hindi 0.004%
+51 Slovenian 0.001%
+n/a Icelandic n/a
+47 Icelandic 0.004%
+68 Bulgarian 0.001%
 48 Bulgarian 0.003%
+78 Icelandic 0.001%
+86 Korean 0.0003%
 102 Hindi 0.0001%
@@
 For Falcon, statistics are estimated based on a sample of the RefinedWeb corpus (roughly 320 million documents).
-Note that for Llama 2 and Falcon, ‘unknown’ indicates texts identified below a predefined confidence threshold (0.5 for Llama 2, 0.65 for our analysis with RefinedWeb).
+Note that for Llama 2 and Falcon, ‘unknown’ indicates texts identified below a predefined confidence threshold of 0.5.
 Touvron et al.
-(2023b) state that this data partially pertain to programming code.
-Target languages for our experiments are highlighted in bold.
-Appendix B Translating AlpacaEval Prompts
+(2023) state that this data partially pertains to programming code.
+Target languages considered in our experiments are highlighted in bold.
+Appendix B Translating Training and Evaluation Data
+As part of our experiments and evaluation, we generate translations of existing datasets.
+All original datasets are licensed under the Apache License 2.0.
+Since translations are generated with GPT-3.5-Turbo, OpenAI’s usage policy applies to the resulting versions, which we make available for future research.
+B.1 AlpacaEval Prompts
 You are a helpful assistant.
@@
 Values in curly braces represent placeholders.
-The value specified for ‘target language’ is English name of the language (e.g., ‘German’) we are translating into, except for Greek and Chinese, where we specify standard modern Greek and Mandarin Chinese respectively.
-To translate AlpacaEval prompts from English into each target language, we query GPT-3.5-Turbo using the template in Figure 8.
-To validate the translated prompts, we manually inspected a sample of the outputs in various languages.
-This inspection revealed that the translations were typically decent, although often included literal translations for metaphorical expressions rather than how a native speaker might express themselves.
+The value specified for ‘target language’ is the English name of the language (e.g., ‘German’) we are translating into, except for Greek and Chinese, where we explicitly specify standard modern Greek and Mandarin Chinese, respectively.
+For the evaluation of single-turn dialogue in non-English languages, we translate AlpacaEval prompts from English into each of our considered target languages using GPT-3.5-Turbo and the template in Figure 9.
+Manual inspection of a sampled subset of these translated prompts in various languages revealed that the translations were typically decent, although often included literal translations for metaphorical expressions rather than how a native speaker might express themselves.
 For instance, English ‘bullet points’ was translated literally into Hindi, rather than an arguably more appropriate phrasing such as ‘important points’.
-Additionally, in languages that distinguish between formal and informal or gendered pronouns (e.g., German, French, Hindi), the formal and male forms are dominant, which may be less representative of how native speakers actually interact with LLM chatbots.
-Appendix C Direct vs. Translated Evaluation for Non-English Chat Responses
-Figure 9: Distribution of helpfulness scores assigned by our LLM judge, GPT-3.5-Turbo, using direct evaluation in the specified target language and after translating target-language responses to English.
-While using a powerful LLM to evaluate the outputs of other models has been shown to achieve reasonable agreement with human judgements in English (Zheng et al., 2023; Chiang and Lee, 2023), it is unclear whether this agreement transfers to all languages under investigation.
-Recent work by Hada et al.
-(2023) has shown that agreement between human and LLM judges tends to be lower for non-English languages, especially in the case of low-resource and non-Latin scripted languages, where the LLM judge tends to be overly optimistic in its assessment.
-However, for certain assessment criteria, such as linguistic acceptability and general content quality, they also confirm that inter-annotator agreement between LLM-based evaluators and humans is inline with that of multiple human annotators.
-To investigate this potential bias, we compared scores assigned by the LLM judge on model-generated responses directly in each non-English target language and on their English translations.
-For each non-English prompt-response pair, we use GPT-3.5-Turbo (gpt-3.5-turbo-1106) and the prompt shown in Figure 8, specifying English as the target language.
-The resulting translated response is then paired with its corresponding AlpacaEval prompt in its original English form and presented to the LLM judge using the prompt in Figure 16.
-Figure 9 shows that the distribution of assigned scores in the direct and translated evaluation settings is very similar for most languages.
-For languages that use non-Latin scripts (e.g., Chinese, Russian, Bulgarian), we observe that GPT-3.5-Turbo tends to assign slightly higher scores more frequently when evaluating directly on the non-English prompt/response pairs.
-This finding agrees with those from Hada et al.
-(2023) and indicates that LLM-based evaluations in non-English languages tend to be slightly overly optimistic and should be considered with caution.
-Nevertheless, we observe that the discrepancy between direct and translated evaluation is relatively minor for the languages considered and leads to negligible differences in the overall average helpfulness score.
-Because of this, and to keep evaluation costs to a minimum, we opt to use the direct evaluation strategy for our experiments.
-Appendix D Translating Non-English Guanaco Training Examples into English
+Additionally, in languages that distinguish between formal, informal, or gendered pronouns (e.g., German, French, etc.), the formal and male forms are dominant.
+While these characteristics may be not truly representative of how native speakers actually interact with LLMs, we consider these to be potentially valid queries for general-purpose chat models.
+B.2 Guanaco Training Examples
 You are a helpful assistant.
@@
 Your translation must also use exact terminology to provide accurate information even for the experts in the related fields.
-Original: {{training_instance}} Translation into {{target_language}}
-Figure 10: Prompt template used to translate Guanaco training instances from English into various target languages.
-In order to investigate the effect of language diversity compared to instruction diversity, we translate a subset of Guanaco’s English training examples into non-English target languages and use these to create MT-based Multi-$i$ instruction-tuning datasets.
+Original: {{training_instance}} Translation into {{target_language}}:
+Figure 10: Prompt template used to translate Guanaco training instances from English into various target languages to investigate the effect of language diversity vs. instruction diversity.
+To investigate the effect of language diversity compared to instruction diversity (§5.1), we translate a subset of Guanaco’s English training examples into non-English target languages and use these to create MT-based Multi-$i$ instruction-tuning datasets.
 By default, speaker roles in Guanaco are denoted with ‘### Human:’ and ‘### Assistant’.
 To ensure that these are never translated and the dialogue structure is maintained, we substitute them with special tokens ‘<S1>’ and ‘<S2>’ and explicitly tell the model to leave these tokens intact (see Figure 10).
-Before training, the special tokens back to their original form.
-Appendix E Results with Falcon 7b
+Before training, we map the special placeholder tokens back to their original form.
+Appendix C Hardware Requirements for Instruction Tuning
+As mentioned in §3, we use LoRA Hu et al.
+(2021) for parameter-efficient finetuning.
+As hyperparameters, we set $R=64$, $\alpha=16$.
+This results in the number of trainable parameters being roughly 2.3% of the original model size for Llama 2 7B, 1.2% for Llama 2 70B and 1.8% for Falcon 7B.
+For 7-billion parameter models, we use two NVIDIA GeForce RTX 3090 with 24GB of memory.
+The time required for each training run is approximately 8 hours.
+For the larger 70-billion parameter model, discussed in §5.2, we use the same hyperparameters and train on four NVIDIA A100 GPUs with 80GB of memory each.
+Here, a single training run takes approximately 20 hours.
+Appendix D MultiSim Datasets
+MultiSim is composed of 34 distinct text simplification data sets aligned at either the document- or sentence-level Ryan et al.
+(2023), covering 13 different languages.
+Due to restrictive licensing on some datasets, only a subset are openly accessible.
+An overview of these datasets is provided in Table 2.
+As can be seen, domain coverage and the amount of available data vary considerably across languages.
+To balance out the number of evaluation instances between the different languages, we randomly sample up to 1,371 complex-simple sentence pairs from all available corpora for each language, except for Slovenian, where we use all 939 available instances.
+While we prioritise sampling from the test and validation splits shared by Ryan et al.
+(2023), we also draw samples from the designated train splits for Portuguese, German, and Slovenian in order to have a sufficiently large evaluation set for our experiments.
+Language Source corpora # instances avail. (test / val / train) # eval. instances
+en ASSET Alva-Manchego et al. (2020) 359 / 100 / 19,000 1,371
+WikiAuto Jiang et al. (2020) 5,002 / 4,988 / 576,126
+ru RSSE Sakhovskiy et al. (2021) 1,083 / 97 / 3,182 1,371
+RuAdapt Encyclopedia Dmitrieva et al. (2021) 839 / 840 / 7,782
+RuAdapt Fairytale Dmitrieva et al. (2021) 31 / 31 / 248
+RuWikiLarge Sakhovskiy et al. (2021) 312 / 678 / 246,978
+de GEOLino Mallinson et al. (2020) 81 / 82 / 958 1,371
+TextComplexityDE Naderi et al. (2019) 26 / 28 / 208
+fr CLEAR Grabar and Cardon (2018) 100 / 294 / 4,196 1,371
+WikiLarge FR Cardon and Grabar (2020) 345 / 878 / 296,402
+pt PorSimples Aluísio and Gasperin (2010) 420 / 420 / 6,290 1,371
+it AdminIT Miliani et al. (2022) 49 / 48 / 588 1,371
+SIMPITIKI Wiki Tonelli et al. (2016) 160 / 146 / 1,436
+PaCCSS-IT Brunato et al. (2016) 1,061 / 1,061 / 60,485
+Teacher Brunato et al. (2015) 17 / 17 / 136
+Terence Brunato et al. (2015) 101 / 102 / 809
+sl SloTS Gorenc and Robnik-Šikonja (2022) 96 / 94 / 749 939
+Table 2: Overview of datasets included in the MultiSim benchmark Ryan et al.
+(2023).
+# instances avail.
+denotes the total number of complex-simple sentence pairs available for each corpus in the MultiSim benchmark.
+# eval.
+instances denotes the number of items sampled from all test, validation, and train splits (where necessary) for language used to evaluate performance on sentence simplification.
+Appendix E Direct vs. Translated Evaluation for Non-English Dialogue Responses
+Figure 11: Distribution of helpfulness scores assigned by our LLM judge, GPT-3.5-Turbo, using direct evaluation in the specified target language and after translating target-language responses to English.
+While using a powerful LLM to evaluate the outputs of other models has been shown to achieve reasonable agreement with human judgements in English Zheng et al.
+(2023); Chiang and Lee (2023), it is unclear whether this agreement transfers to all languages under investigation.
+Recent work by Hada et al.
+(2024) has shown that agreement between human and LLM judges tends to be lower for non-English languages, especially in the case of low-resource and non-Latin scripted languages, where the LLM judge tends to be overly optimistic in its assessment.
+However, for certain assessment criteria, such as linguistic acceptability and general content quality, they also confirm that inter-annotator agreement between LLM-based evaluators and humans is in line with that of multiple human annotators.
+To investigate this potential bias, we compared scores assigned by the LLM judge on model-generated responses directly in each non-English target language and on their English translations.
+For each non-English prompt-response pair, we translate it into English using GPT-3.5-Turbo (gpt-3.5-turbo-1106) and the prompt shown in Figure 9.
+We pair the resulting translated responses with their corresponding English AlpacaEval prompts and repeat our LLM judge evaluation.
+Figure 11 shows that the distribution of assigned scores in the direct and translated evaluation settings is very similar for most languages.
+For languages that use non-Latin scripts (e.g., Chinese, Russian, Korean, etc.), we observe that the LLM judge tends to assign slightly higher scores more frequently when evaluating directly on the non-English prompt/response pairs.
+This finding agrees with those from Hada et al.
+(2024) and indicates that LLM-based evaluations in non-English languages can be overly optimistic and should be considered with caution.
+Nevertheless, we observe that the discrepancy between direct and translated evaluation is relatively minor for the languages considered and leads to negligible differences in the overall average helpfulness score.
+Based on these results, we opt to use the direct evaluation strategy for our experiments, which has the added benefit of avoiding the introduction of potential translation errors and keeping the cost of evaluation to a minimum.
+Appendix F Results with Falcon 7b
 In order to assess whether our findings generalise to other LLMs, we repeat our experiments using Falcon 7b Almazrouei et al. (2023).
-E.1 Chat
-Figure 11 shows the helpfulness scores assigned by our LLM judge for Falcon 7b given incremental multilingual instruction tuning across all target languages.
-Similarly to our results with Llama 2 7b (see Figure 2), cross-lingual transfer is elicited after finetuning with relatively few languages, and no additional gains observed when including more than three languages.
-That said, Falcon 7b appears to show strong performance on French, even without multilingual finetuning, indicating that, despite being an English-centric model, it has strong capabilities in French out of the box.
-For Spanish, German and Chinese, performance is comparable to that of Llama 2 7b.
+F.1 Single-turn Dialogue
+Figure 12 shows the helpfulness scores assigned by the LLM judge for Falcon 7b given incremental multilingual instruction tuning across all target languages.
+Similarly to our results with Llama 2 7b (cf.
+Figure 2), cross-lingual transfer can be elicited after finetuning with relatively few languages, and no additional gains observed when including more than three languages.
+Interestingly, Falcon 7b appears to show strong performance on French, even without multilingual finetuning, indicating that, despite being an English-centric model, it has strong capabilities in French out of the box.
+For Spanish, German, and Chinese, performance is comparable to that of Llama 2 7b.
 However, for all other languages, responses are often ranked least helpful, indicating that Falcon 7b’s multilingual capabilities are limited strictly to major European languages and Chinese.
-E.2 Extractive Question Answering
+F.2 Sentence Simplification
+Figure 13 depicts Falcon 7b’s performance on the sentence simplification task given multilingual instruction tuning.
+In general, the results emulate those seen with Llama 2 7b.
+Performance on English remains uniform as multilinguality increases, while for non-English target languages, the largest jump in performance happens when moving from monolingual to bilingual instruction tuning, and plateauing again with just three languages.
+The notable exceptions here are French and Italian, for which Falcon 7b appears to provide greater support even in the monolingual instruction tuning setting, echoing the results from the single-turn dialogue task.
+F.3 Extractive Question Answering
 Figure 14 shows the results of Falcon 7b on XQuAD.
-While performance is generally lower than that of Llama 2 7b on this task (see Figure 3), we observe a similar, albeit weaker, effect of multilingual finetuning within the supported languages (Spanish, German, Chinese).
-E.3 Commonsense Reasoning
+While performance is generally lower than that of Llama 2 7b on this task (cf.
+Figure 4), we observe a similar, albeit weaker, effect of multilingual finetuning within the supported languages (es, de, zh).
+F.4 Commonsense Reasoning
 Figure 15 shows the results of Falcon 7b on X-CSQA.
-Strikingly, in contrast to the results achieved with Llama 2 7b (Figure 4), Falcon 7b fails to score above random performance across all target languages.
+Strikingly, in contrast to the results achieved with Llama 2 7b (cf.
+Figure 5), Falcon 7b fails to score above random performance across all target languages.
 Regarding the effect of multilingual instruction tuning, we again see that it fails to deliver any performance improvements on this highly structured task.
-E.4 XNLI
+Degradation on Russian and Vietnamese, is likely due to the model consistently failing to follow instructions and reproducing part of the question as the response.
+F.5 XNLI
 Figure 16 shows the results of Falcon 7b on XNLI.
-Similar to the results attained with Llama 2 7b (see Figure 5), we observe no significant differences in performance given different degrees of multilingual instruction tuning.
-E.5 Discussion
-Overall, these experiments indicate that Falcon’s multilingual capabilities are considerably narrower than Llama 2’s.
-We hypothesise that this may be a result of more stringent filtering of web-scraped pretraining data Penedo et al.
-(2023b), possibly resulting in less accidental contamination.
-Interestingly though, according to our estimated statistics based on the sample of the RefinedWeb corpus (Table 1), which makes up a large portion of Falcon’s pretraining data, Russian appears as frequently as Chinese and Swedish.
-However, we find that Russian texts are, on average, three times longer when tokenized with Falcon’s tokenizer compared to Llama 2’s.
-This indicates that, despite having a much larger vocabulary (65k for Falcon vs.
-32k for Llama 2), Cyrillic-scripted languages such as Russian and Bulgarian are outside of the model’s intended domain and performance on these languages is inherently limited.
-Figure 11: Average helpfulness of chat responses from Falcon 7b with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
+Similar to the results attained with Llama 2 7b (cf.
+Figure 6), we observe no significant differences in performance given different degrees of multilingual instruction tuning.
+F.6 Discussion
+These additional experiments using Falcon 7b provide further support for our main findings: instruction tuning with as few as two languages significantly improves cross-lingual generalisation, enhancing performance in open-ended single-turn dialogue for some non-English languages.
+However, strikingly, multilingual capabilities of Falcon are considerably narrower than that of Llama 2.
+This is also suggested by the statistics provided in Table 1, which show a higher concentration of English and much lower proportions of non-English languages compared with both Llama 2 and GPT-3.
+We suspect that this may be a result of stringent filtering of web-scraped pretraining data performed in producing the RefinedWeb corpus used to train the Falcon models Penedo et al.
+(2023), which not only reduces the risk of potentially accidental contamination Blevins and Zettlemoyer (2022) but also language coverage.
+That said, the RefinedWeb corpus comprises only part of Falcon’s actual training corpus.
+Almazrouei et al.
+(2023) note that the final corpus also contains additional data drawn from curated sources including a European-focused (multilingual) Common-Crawl dataset, which could explain the strong performance on some major European languages.
+In light of this, we suspect that additional Chinese data is also included in Falcon’s training corpus, allowing for relatively strong performance in Chinese on the single-turn dialogue task.
+Figure 12: Average helpfulness of single-turn dialogue responses from Falcon 7b given incremental multilingual instruction tuning.
+Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 13: SARI weighted by IO language agreement for sentence simplification with Falcon 7b given incremental multilingual instruction tuning.
+Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx) Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 14: XQuAD results for Falcon 7b given incremental multilingual instruction tuning.
+Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx).
+Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 15: X-CSQA results for Falcon 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 16: XNLI results for Falcon 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Appendix G Results with Llama 3 8b
+In §5.2 we observed that scaling the base model from 7b to 70b parameters reduces the gap in performance between English and non-English target languages with regard to the single-turn dialogue task.
+In this section, we evaluate Llama 3 8b MetaAI (2024), which builds on Llama 2 by scaling up the pretraining data from 2 trillion tokens to 15 trillion tokens.
+In Addition to a much larger pretraining corpus, Llama 3 is trained with a vocabulary of 128k tokens (vs.
+32k for Llama 2), bringing it much closer to traditional multilingual models which typically contain roughly 250k tokens Le Scao et al.
+(2023); Xue et al.
+(2021); Conneau et al.
+(2020).
+While the exact language distribution of the model’s pretraining data is not publicly known, it is reasonable to expect that these enhancements could lead to improvements in the model’s non-English capabilities.
+Figure 17: Single-turn dialogue performance of Llama 3 8b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
+Figure 17 shows the performance of Llama 3 8b, which closely resembles that of Llama 2 70b across most non-English languages, even outperforming it in languages like Vietnamese, Greek, and Hindi.
+Most notably, we observe that our main findings still hold: multilingual instruction tuning with as few as two languages is both necessary and sufficient to elicit cross-lingual generalisation.
+Again, performance typically plateaus thereafter.
+Given the improved performance on languages like Greek and Hindi that are low-resource under Llama 2’s pretraining data distribution, we suspect that their representation in the pretraining data for Llama 3 is substantially higher.
+Still, performance on Icelandic suggests that this language may remain in the low-resource category under this model.
+Appendix H Ablation Experiments
+In Section 3.2, we constructed Multi-$i$ instruction tuning datasets of the same size by replacing a fixed number of English examples with examples from a new language.
+As a result, the proportion of non-English examples used for finetuning increases along with the number of languages.
+Table 3 shows the exact makeup of each these datasets and how they relate to one another.
+To rule out potential confounding factors between the proportion of non-English examples and the diversity of languages, we conduct ablation experiments in which we keep one variable fixed while incrementing the other.
+Dataset Languages % NE Total
+Mono en 0.00% 3,200
+Multi-2 en, es 6.25% 3,200
+Multi-3 en, es, ru 12.50% 3,200
+Multi-4 en, es, ru, de 18.75% 3,200
+Multi-5 en, es, ru, de, zh 25.00% 3,200
+Multi-6 en, es, ru, de, zh, fr 31.25% 3,200
+Table 3: Makeup of the incremental multilingual instruction-tuning datasets.
+[…]
+(b) Impact of incrementing the total number of non-English examples while keeping the number of languages fixed at three (English, Spanish, and Russian).
+Here, Russian is seen during finetuning, while German and Chinese are in the 0-shot setting.
+Figure 18: Average helpfulness of single-turn dialogue responses from Llama 2 7b given multilingual instruction tuning on a fixed budget (limiting either the number of non-English instances 18(a) or the number of languages 18(b)).
 Error bars show a confidence interval of 95%.
-Figure 12: X-QuAD results for Falcon 7b with incremental multilingual instruction tuning.
-Results are shown for both multilingual prompting (en:xx) and monolingual prompting (xx:xx).
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
-Figure 13: X-CSQA results for Falcon 7b with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
-Figure 14: XNLI results for Falcon 7b with incremental multilingual instruction tuning.
-Striped bars indicate that the target language is included in the finetuning data.
-Error bars show a confidence interval of 95%.
+Figure 18(a) shows that when using a fixed budget of 400 non-English examples (which equates to 12.5% of the finetuning data) and incrementing only the number of languages within this budget, performance increases consistently as each new language is added, closely reflecting the results of our main experiments.
+In contrast, Figure 18(b) shows that when training with a fixed number languages (e.g., Multi-2) and incrementing the number of non-English examples (in this case, Spanish), performance tends to remain lower and generally shows less improvement.
+These results underscore our main findings discussed in §4.1 and §6.
 Figure 19: Comparison of Llama 2 7b with multilingual instruction tuning using native (Native NE) and translated non-English (MT NE) examples for additional target languages not shown in Figure 7.
@@
 You are an expert language evaluator.
-You are evaluating a response that has been submitted for a particular task, using a specific set of standards.
-Below is the data: [BEGIN DATA] *** [Task]: {{task}} *** [Submission]: {{submission}} *** [Criterion]: helpfulness: ‘‘1’’: ‘‘Not helpful - The generated text is completely irrelevant, unclear, or incomplete.
-It does not provide any useful information to the user.’’ ‘‘2’’: ‘‘Somewhat helpful - The generated text has some relevance to the user’s question, but it may be unclear or incomplete.
-It provides only partial information, or the information provided may not be useful for the user’s needs.’’ ‘‘3’’: ‘‘Moderately helpful - The generated text is relevant to the user’s question, and it provides a clear and complete answer.
-However, it may lack detail or explanation that would be helpful for the user.’’ ‘‘4’’: ‘‘Helpful - The generated text is quite relevant to the user’s question, and it provides a clear, complete, and detailed answer.
+You are evaluating a response that has been submitted for a particular task, using a specific set of standards. Below is the data:
+[BEGIN DATA]
+[Task]: {{task}}
+[Submission]: {{submission}}
+[Criterion]: helpfulness:
+‘‘1’’: ‘‘Not helpful - The generated text is completely irrelevant, unclear, or incomplete. It does not provide any useful information to the user.’’
+‘‘2’’: ‘‘Somewhat helpful - The generated text has some relevance to the user’s question, but it may be unclear or incomplete.
+It provides only partial information, or the information provided may not be useful for the user’s needs.’’
+‘‘3’’: ‘‘Moderately helpful - The generated text is relevant to the user’s question, and it provides a clear and complete answer.
+However, it may lack detail or explanation that would be helpful for the user.’’
+‘‘4’’: ‘‘Helpful - The generated text is quite relevant to the user’s question, and it provides a clear, complete, and detailed answer.
 It offers additional information or explanations that are useful for the user.
-However, some of the points of the response are somewhat repetitive or could be combined for greater clarity and concision.’’ ‘‘5’’: ‘‘Very helpful - The generated text is highly relevant to the user’s question, and it provides a clear, complete, and detailed answer.
+However, some of the points of the response are somewhat repetitive or could be combined for greater clarity and concision.’’
+‘‘5’’: ‘‘Very helpful - The generated text is highly relevant to the user’s question, and it provides a clear, complete, and detailed answer.
 It offers additional information, explanations, or analogies that are not only useful but also insightful and valuable to the user.
-However, the structured of the response is not well-organized and there is no clear progression or logical sequence of different points in the response.’’ ‘‘6’’: ‘‘Highly helpful - The generated text provides a clear, complete, and detailed answer.
+However, the structured of the response is not well-organized and there is no clear progression or logical sequence of different points in the response.’’
+‘‘6’’: ‘‘Highly helpful - The generated text provides a clear, complete, and detailed answer.
 It offers additional information or explanations that are not only useful but also insightful and valuable to the user.
-The response is also in a logical and easy-to-follow manner by explicitly using headings, bullet points, or numbered lists to break up the information and make it easier to read.’’ *** [END DATA] Does the submission meet the criterion?
+The response is also in a logical and easy-to-follow manner by explicitly using headings, bullet points, or numbered lists to break up the information and make it easier to read.’’
+[END DATA]
+Does the submission meet the criterion?
 First, write out in a step by step manner your reasoning about the criterion to be sure that your conclusion is correct.
 Avoid simply stating the correct answers at the outset.
-Provide your response as a RFC8259 compliant JSON following this schema: {‘‘reasoning’’: str, ‘‘score’’: int}
-Figure 16: Prompt template used to query the LLM judge in order to assess the helpfulness of an open-ended chat response.
+Provide your response as a RFC8259 compliant JSON following this schema:
+{‘‘reasoning’’: str, ‘‘score’’: int}
+Figure 20: Prompt template used to query the LLM judge in order to assess the helpfulness of a single-turn dialogue response.
 This prompt closely follows the one used by Zhou et al.
 (2023) but extends it by specifying the return value as a valid JSON-like object.
-Prompt:
-### Human: Answer the question from the given passage. Your answer should be directly extracted from the passage, and it should be a single entity, name, or number, not a sentence.
+Prompt: ### Human: Please rewrite the following complex sentence in order to make it easier to understand.
+You can do so by replacing complex words with simpler synonyms (i.e.
+paraphrasing), deleting unimportant information (i.e.
+compression), and/or splitting a long complex sentence into several simpler ones.
+The final simplified sentence needs to be grammatical, fluent, and retain the main ideas of its original counterpart without altering its meaning.
+Complex sentence: Sie gelten als eiskalte Fress-Maschinen.
+### Assistant:
+Reference: Sie gelten als Fress-Maschinen.
+System: They are seen as cold-blooded eaters.
+Table 4: An example of the 0-shot sentence simplification task.
+Monospaced font represents the prompt template used, adapted from instructions provided to crowdworkers for the creation of the ASSET dataset Alva-Manchego et al.
+(2020).
+This example shows the cross-lingual prompt strategy (en:xx), where the task instruction is provided in English but the source sentence is in German.
+Ideally, the system language should match the source sentence, which it fails to do in this case.
+Prompt: ### Human: Answer the question from the given passage. Your answer should be directly extracted from the passage, and it should be a single entity, name, or number, not a sentence.
 Passage:
@@
 ### Assistant: Based on the passage, the answer to the question is ‘‘
-Reference:
-zwei
-System:
-two".
+Reference: zwei
+System: two".
 In the passage, it is stated that Josh Norman “erzielte vier Interceptions, von denen zwei zu Touchdowns für sein Team wurden.” This means that Norman intercepted four passes during the season, and two of those interceptions were returned for touchdowns.
-Table 2: Example input prompt, reference and model response for the XQuAD task.
+Table 5: Example input prompt, reference, and model response for the 0-shot extractive QA task using XQuAD.
 Monospaced font represents the prompt template used for each test instance.
-Here, we show the multilingual prompting strategy (en:xx), where the task instruction and response prefix is provided English, while the context passage and question are in the target language (in this case, German).
+Here, we show the cross-lingual prompting strategy (en:xx), where the task instruction and response prefix is provided English, while the context passage and question are in the target language (in this case, German).
 A correct answer in this setting requires the model to continue the response by switching into the target language, which it fails to do in this particular example.
-Prompt:
-### Human: Dir wird eine Frage vorgelegt, die mehrere Antwortmöglichkeiten hat.
+Prompt: ### Human: Dir wird eine Frage vorgelegt, die mehrere Antwortmöglichkeiten hat.
 Wähle die am besten geeignete Option aus ‘‘A’’, ‘‘B’’, ‘‘C’’, ‘‘D’’ oder ‘‘E’’ basierend auf deinem gesunden Menschenverstand.
@@
 ### Assistant: ‘‘
-Reference:
-E
-System:
-E”
-Table 3: Example of the X-CSQA task.
+Reference: E
+System: E”
+Table 6: Example of the commonsense reasoning task using X-CSQA.
 Monospaced font represents the prompt template used for each test instance.
@@
 A correct answer requires the model to output the letter corresponding to the ground-truth answer.
+XNLI prompt template: {Sentence 1}, right? {label}, {Sentence 2}
+Label Prompt query NLL
+Entailment In measuring effectiveness, perfection is unattainable., right? Yes, You can never be perfect. -113.94
+Neutral In measuring effectiveness, perfection is unattainable., right? Also, You can never be perfect. -116.79
+Contradiction In measuring effectiveness, perfection is unattainable., right? No, You can never be perfect. -114.01
+Table 7: Example prompt queries for XNLI given the predefined template used in the LM Evaluation Harness Gao et al.
+(2023) and the three possible labels.
+Underlined words are the language-specific connectors corresponding to the three possible labels.
+The final answer is the sequence with the highest negative log-likelihood (NLL) according to the model.
+In this case, Llama 2 7b Mono correctly identifies the relationship between the two sentences as entailment.